### v1

# Deep polytopic autoencoders for low-dimensional linear parameter-varying approximations and nonlinear feedback design

Abstract
Polytopic autoencoders provide low-dimensional parametrizations of states in a polytope. For nonlinear PDEs, this is readily applied to low-dimensional linear parameter-varying (LPV) approximations as they have been exploited for efficient nonlinear controller design via series expansions of the solution to the state-dependent Riccati equation. In this work, we develop a polytopic autoencoder for control applications and show how it outperforms standard linear approaches in view of LPV approximations of nonlinear systems and how the particular architecture enables higher order series expansions at little extra computational effort. We illustrate the properties and potentials of this approach to computational nonlinear controller design for large-scale systems with a thorough numerical study.
[…]
1 Introduction
The design of controllers for nonlinear high-dimensional systems is a challenging task and there exists no generally applicable and computationally feasible approach. Here, we consider input-affine systems of the form
$\dot{v}(t)=f(v(t))+Bu(t),\quad y(t)=Cv(t),$ (1)
where for time $t>0$, $v(t)\in\mathbb{R}^{n}$ denotes the state, $u(t)\in\mathbb{R}^{m}$ and $y(t)\in\mathbb{R}^{\ell}$ denote the system’s input and output, $f\colon\mathbb{R}^{n}\to\mathbb{R}^{n}$ is a possibly nonlinear function, and $B$ and $C$ are linear input and output operators.
For the control of systems like eq. 1, methods like backstepping [17], feedback linearization [19, Chap. 5.3] and sliding mode control [12] require structural assumptions whereas general approaches based on the Hamilton-Jacobi-Bellman (HJB) equations are only feasible for very moderate system sizes; see, e.g., [8]. The commonly employed state-dependent Riccati equation (SDRE) approximation to the HJB (see, e.g., [10] for a survey) lifts a significant portion of the computational complexity. Nonetheless, this approach became available for large-scale systems only in recently due to further simplifications; see [6] for a linear update scheme and [1] for approximations through series expansions.
In this work, we expand on the series expansion of the SDRE combined with the approximation of the nonlinear system through a low-dimensional linear parameter-varying (LPV) system. This allows us to provide a generic algorithm to nonlinear controller design; see our previous works [16, 11] and related discussions in [1, 2]. In particular, the results in [16] indicate a measurable improvement of the nonlinear feedback design over a related, well-tuned and robust LQR controller but left space for further performance improvements, for example, through
  1. 1.
    higher order expansions of the nonlinear feedback or
  2. 2.
    nonlinear parametrizations of the velocity states.
Both topics are linked to the question if a nonlinear parametrization can provide better low-dimensional approximations with fewer degrees of freedom and thus mitigating the natural limit, which is commonly expressed through the Kolmogorov $n$-width of any linear model order reduction scheme – like proper orthogonal decomposition (POD) which is a linearly optimal projection scheme; [13]. There have been many efforts and positive results concerning the use of nonlinear model order reduction approaches using, e.g., polynomial manifold maps [9], learning techniques [13], or the selection of local bases [3].
To consider higher order expansions of a nonlinear feedback law, a very low-dimensional parametrization is crucial. Among others, as we will lay out, the numbers of coefficient that need to be precomputed through large-scale matrix equations equals $\sum_{k=0}^{p}\binom{r+k-1}{k}$, where $r$ is the dimension of the parametrization and $p$ is the order of the expansion. For example, for $r=10$ and a first-order expansion, the solve of $1+10=11$ matrix equations was required, cp. [16], whereas a second or third-order expansion would require $66$ or $286$ solves. However, for a parametrization of size $r=5$, only $6$, $21$, or $56$ matrix equation solves are needed for expansions of first, second and third order, respectively.
In this work, we rely on our recent development of polytopic autoencoders [14] that combines learning approaches with the selection of local bases in terms of polytope vertices. The concept of reconstructing the states in a polytope provides certain advantages over the reconstruction in a linear space. Generally, the observed states of a dynamical system are described in a bounded set rather than in an unbounded space. With the polytopic approach, the reconstruction is confined to a region where the states have been observed. This confinement stands against the freedom to explore the full state space, for controller design, however, it can be beneficial if tailored towards target regimes and since the resulting series expressions will be bounded by design.
The underlying ideas and major findings of this paper can be summarized as follows: We recall in Section 2 that a general nonlinear control system can be approximated by a low-dimensional LPV system, which can be exploited for controller design in particular if the parameter dependency itself is linear. In view of providing a performant but maximally low-dimensional parametrization, we resort to clustered polytopic autoencoders that encode the states of a dynamical system in a polytope and decode them with locally adapted bases; see Section 3. An interesting result in view of controller design is that this nonlinear decoding has a Taylor expansion with vanishing second-order terms (Lemma 1). For the actual controller design in Section 4, we consider series expansions of state-dependent Riccati equation as it has been proposed and exercised before. Here, we extend the formulas to consider nonlinear parameter dependencies and higher order terms in the expansion. Finally, we report numerical experiments that illustrate the performance improvement of the polytopic autoencoder over standard approximation schemes, that explore the different levels of approximation to the underlying nonlinear SDRE feedback design, and that investigate the gain in controller performance for a commonly considered flow control problem; see Section 6.
2 Low-dimensional LPV approximations
For nonlinear systems like eq. 1, we consider quasi-LPV approximations of the form
$\dot{v}(t)=A\bigl{(}\rho(v(t))\bigr{)}\,v(t)+Bu(t),\quad y(t)=Cv(t).$ (2)
Such systems can be derived from a state-dependent coefficient representation such as $f(v)=\widetilde{A}(v)v$ and a parametrization of the state $v(t)\approx\tilde{v}(t)=\nu(\mu(v(t))$, where $\mu$ is an encoder and $\nu$ is a decoder. In this case, we set $\rho=\mu(v)$ and $A(\rho):=\widetilde{A}(\nu(\rho))$. Note that the parameter depends on the state, hence the term quasi-LPV.
Remark 1.
In the example application we consider later, namely flows, we will work with coefficients $\widetilde{A}$ that are linear in the variable so that linear structures in the decoding $\rho\to\tilde{v}$ will lead to linear parameter dependencies as they have been proven beneficial for controller design; see, e.g., [4].
3 Polytopic parametrizations in clusters for controller design
As opposed to linear model order reduction approaches, where the state is parametrized in a linear subspace, in polytopic autoencoders (PAEs), the state is represented through coordinates with respect to vertices of an $R$-dimensional polytope. As a polytope is bounded, without loss of generality, the vertices $\tilde{v}_{k}$ can be scaled and shifted such that the coordinates $\theta(t)\in\mathbb{R}^{R}$ that represent the reconstruction $\tilde{v}(t)$ in the polytope will be positive and sum up to one, i.e.,
$\tilde{v}(t)=\sum_{k=1}^{R}\theta_{k}(t)\tilde{v}_{k},\quad\text{with}\leavevmode\nobreak\ \theta(t)\geq 0,\quad\sum_{k=1}^{R}\theta_{k}(t)=1.$
This design is not capable of exactly parametrizing a target value $v^{*}=0$ unless the support vectors $\tilde{v}_{k}$ come with a specific linear dependency. To explicitly enable the reconstruction of the zero target value and to ensure that the series expansion of the SDRE is about $\theta(\tilde{v})=0$, we include $\tilde{v}_{0}=0$ in the reconstruction basis of the training, where we impose the positivity and summation constraint on the $\theta$ variable extended by $\theta_{0}$, i.e.,
$\theta_{0}(t)+\sum_{k=1}^{R}\theta_{k}(t)=1.$ (3)
This way, the tuple $(1,\theta(t))=(1,0)\in\mathbb{R}^{R+1}$ will be a feasible value in the extended parameter space that represents the target value.
Following previous results that showed how clustering, i.e., how the choice of local bases can improve the reconstruction, we combine the polytopic parametrization with smooth clustering networks as we have developed it in [14]. In short, we will use coordinates $\theta=c(\rho)\otimes\rho$, where $\rho\in\mathbb{R}^{r}$ denotes the actual latent variable that encodes $v$ and where $c$ is a smooth clustering network that blends between $q$ sets of local bases for the reconstruction.
Taking into consideration a zero target value, we propose a polytopic autoencoder that consists of three components:
  1. 1.
    A nonlinear encoder $\mu\colon\mathbb{R}^{n}\to\mathbb{R}^{r+1}$ so that
    $(\rho_{0},\rho)=\mu(v)=:\operatorname{softmax}(\operatorname{F_{e}}(v)+e_{1}),$
    with a scalar function $\rho_{0}$ and $\rho(t)\in\mathbb{R}^{r}$, where the function $\operatorname{F_{e}}\colon\mathbb{R}^{n}\to\mathbb{R}^{r+1}$ is a trainable nonlinear convolutional neural network with $\operatorname{F_{e}}(0)=0$ and where $e_{1}\in\mathbb{R}^{r+1}$ is the first unit vector. Since $\operatorname{F_{e}}(0)=0$ and with the softmax function
    $\operatorname{softmax}\colon x\to\left(\tfrac{x_{i}\tanh(ax_{i})}{\sum_{j=1}^{r+1}x_{j}\tanh(ax_{j})}\right)_{i}^{r+1},$
    for a parameter $a>0$ (that we will set to $a=10$ in the experiments), we ensure that the output $(\rho_{0}(t),\rho(t))$ satisfies
    $(\rho_{0}(t),\rho_{k}(t))\geq 0\leavevmode\nobreak\ \leavevmode\nobreak\ \text{and}\leavevmode\nobreak\ \leavevmode\nobreak\ \rho_{0}(t)+\sum_{k=1}^{r}\rho_{k}(t)=1;$
    in particular $\mu(0)=(1,0)$.
  2. 2.
    A smooth clustering network $c\colon\mathbb{R}^{r}\to\mathbb{R}^{q}$ so that
    $c(\rho)=\operatorname{softmax}(\operatorname{F_{c}}(\rho)+\hat{e}_{1}),$ (4)
    where $\operatorname{F_{c}}$ is a trainable neural network with $\operatorname{F_{c}}(0)=0$ and where $\hat{e}_{1}$ is the first unit vector in $\mathbb{R}^{q}$. By design, for the output $c(\rho)$, we have that
    $c(\rho(t))\geq 0\quad\text{and}\quad\sum_{i=1}^{q}c(\rho(t))_{i}=1$
    hold, so that $(\theta_{0},\theta):=(\rho_{0},c(\rho)\otimes\rho)$ satisfies the positivity and summation conditions eq. 3.
  3. 3.
    A trainable linear map $W\in\mathbb{R}^{n\times qr}$ that realizes the reconstruction as
    $\tilde{v}=W\big{(}c(\rho)\otimes\rho\big{)}.$ (5)
We lay out the design and motivation in the following remarks:
  • The property that $\operatorname{F_{e}}(0)=0$ and $\operatorname{F_{c}}(0)=0$ is achieved by using (convolutional) nonlinear layers without bias terms and with activation functions that cross the origin.
  • The use of the extra $\rho_{0}$ variable ensures that $\rho$ is zero at the origin despite being valid coordinates in a polytope.
  • The design of the clustering network ensures $c(0)=e_{1}$ so that values of $\rho$ close to zero are associated with the first cluster and that $c(\mu(0))\otimes\mu(0)=(1,0)\in\mathbb{R}^{qr}$ represents the target vector $v^{*}=0$ exactly.
  • With using the Kronecker product $c(\rho)\otimes\rho$, the $k$-th $r$ columns of $W$ denote the supporting basis for reconstruction in the $k$-th cluster; see [14].
Lemma 1.
With $\operatorname{softmax}\colon x\to\left(\tfrac{x_{i}\tanh(ax_{i})}{\sum_{j=1}^{q}x_{j}\tanh(ax_{j})}\right)_{i}^{q}$, it holds that the Jacobian $c^{\prime}$ of the clustering network $c$ as defined in eq. 4 is zero at $\rho=0$. In particular, the second-order Taylor expansion of $\rho\to A(\rho)$ coincides with the first-order expansion and reads
$A(\rho)\approx A(0)+\sum_{j}^{r}\rho_{j}\widetilde{A}(w_{j}),$ (6)
where $w_{j}$ are the first $r$ columns of $W\in\mathbb{R}^{n\times qr}$ as in eq. 5.
Proof.
By direct computations we find $c^{\prime}(\operatorname{F_{c}}(\rho)+\hat{e}_{1})\big{\rvert}_{\rho=0}=c^{\prime}(\hat{e}_{1})=0\in\mathbb{R}^{q\times r}$. Since the coefficient map $c(\rho)\otimes\rho\to\widetilde{A}(\tilde{v}(\rho))$ is linear, we can reduce the argument to the nonlinear part $\rho\to c(\rho)\otimes\rho$. With $\rho=0$, from
$c(\rho+h)\otimes(\rho+h)-c(\rho)\otimes\rho=c(0+h)\otimes h,$
we derive the first and second-order expansion terms of the map $\rho\to c(\rho)\otimes\rho$ in the increment $h$ at $\rho=0$ as
$c(0+h)\otimes h=\left(c(0)+c^{\prime}(0)h+o(\lvert h\rvert^{2})\right)\otimes h$
and as
$c(\rho)\otimes\rho\approx c(0)\otimes\rho+(c^{\prime}(0)\rho)\otimes\rho=c(0)\otimes\rho+(c^{\prime}(0)\otimes I)(\rho\otimes\rho),$
where $c^{\prime}(0)\in\mathbb{R}^{q\times r}$ is the Jacobian of $c$ at $\rho=0$. With $c^{\prime}(0)=0$, the second-order terms vanish so that with $c(0)=\hat{e}_{1}$, the first (and second-order) approximation to the polytopic reconstruction reads
$\tilde{v}(\rho)\approx W\left(\hat{e}_{1}\otimes\rho\right)=\sum_{j=1}^{r}\rho_{j}w_{j}.$ (7)
Therefore, formula eq. 6 follows from the linearity of $\widetilde{A}$. ∎
4 Higher order series expansions of LPV approximations in SDRE feedback design
For the quasi-LPV system eq. 2, the nonlinear SDRE feedback design defines the input $u$ as
$u(t)=-B^{\mkern-1.5mu\mathsf{T}}P(\rho(v(t)))v(t),$ (8)
where $P(\rho)$ solves the SDRE
$A(\rho)^{\mkern-1.5mu\mathsf{T}}P(\rho)+P(\rho)A(\rho)-P(\rho)BB^{\mkern-1.5mu\mathsf{T}}P(\rho)=-C^{\mkern-1.5mu\mathsf{T}}C.$ (9)
Since the repeated solve of eq. 9 is costly, series expansions of the feedback law eq. 8 have been proposed.
We briefly derive the formulas for higher order expansions with nonlinear parameter dependencies in $A$. For linear parameter dependencies, the first-order approximation has been developed in [5] for a single parameter dependency and has been extended in [1] to the multivariate case.
If the state dependency of the SDRE solution is parametrized through $\rho$, the multivariate Taylor expansion of $P$ about $\rho_{0}=0$ up to order $p$ reads
$P(\rho)\approx P_{0}+\sum_{1\leq\lvert\alpha\rvert\leq p}\rho^{(\alpha)}P_{\alpha},$
where $\alpha=(\alpha_{1},\ldots,\alpha_{r})\in\mathbb{N}^{r}$ is a multiindex with $\lvert\alpha\rvert:=\sum_{i=1}^{r}\alpha_{i}$, where $\rho^{(\alpha)}:=\rho_{1}^{\alpha_{1}}\rho_{2}^{\alpha_{2}}\cdots\rho_{r}^{\alpha_{r}}$, and where, importantly, $P_{\alpha}$ are constant; see [16] for more details.
For nonlinear parameter dependencies, we will consider a truncated series expansion of $A$:
$A(\rho)\approx A_{0}+\sum_{1\leq\lvert\alpha\rvert\leq p}\rho^{(\alpha)}A_{\alpha},$ (10)
where $A_{\alpha}$ are constant matrices; see eq. 6 that describes the first and second order expansion in the considered case with the structure of the nonlinear decoding as $\rho\to W(c(\rho)\otimes\rho)$.
We insert the expansions of $A$ and $P$ into the SDRE eq. 9 and solve for the coefficients $P_{\alpha}$ by matching the individual powers of $\rho$. For $\alpha=0$, we obtain the standard Riccati equation
$A_{0}^{\mkern-1.5mu\mathsf{T}}P_{0}+P_{0}A_{0}-P_{0}BB^{\mkern-1.5mu\mathsf{T}}P_{0}=-C^{\mkern-1.5mu\mathsf{T}}C,$ (11)
whereas for $\lvert\alpha\rvert=1$, the coefficients $P_{\alpha}$ are obtained from
$\begin{split}P_{0}\bigl{(}\sum_{\lvert\alpha\rvert=1}\rho^{(\alpha)}A_{\alpha}\bigr{)}+\bigl{(}\sum_{\lvert\alpha\rvert=1}\rho^{(\alpha)}A_{\alpha}\bigr{)}^{\mkern-1.5mu\mathsf{T}}P_{0}\\ {}-{}\sum_{\lvert\alpha\rvert=1}\rho^{(\alpha)}P_{\alpha}BB^{\mkern-1.5mu\mathsf{T}}P_{0}-P_{0}BB^{\mkern-1.5mu\mathsf{T}}\sum_{\lvert\alpha\rvert=1}\rho^{(\alpha)}P_{\alpha}=0\end{split}$
for arbitrary values of $\rho$. This amounts for the $P_{\alpha}$ to be given via the Lyapunov equations
$\left(A_{0}-BB^{\mkern-1.5mu\mathsf{T}}P_{0}\right)^{\mkern-1.5mu\mathsf{T}}P_{\alpha}+P_{\alpha}\left(A_{0}-BB^{\mkern-1.5mu\mathsf{T}}P_{0}\right)$
$=-A_{\alpha}^{\mkern-1.5mu\mathsf{T}}P_{0}-P_{0}A_{\alpha},$ (12)
for the $r$ multiindices in $\alpha\in\mathbb{N}^{r}$ for which $\lvert\alpha\rvert=1$; see also [1, 16] where the same formulas have been derived. Similarly, we may collect the terms in which $\rho$ appears quadratic. Let ${\alpha^{*}}$ be a multiindex with $\lvert{\alpha^{*}}\rvert=2$ and let ${\beta^{*}}$ and ${\delta^{*}}$ be those two multiindices with $\lvert{\beta^{*}}\rvert=\lvert{\delta^{*}}\rvert=1$ that satisfy ${\beta^{*}}+{\delta^{*}}={\alpha^{*}}$. Then, the coefficient $P_{{\alpha^{*}}}$ that cancels the corresponding contributions in the equation is defined through the Lyapunov equation
$(A_{0}-BB^{\mkern-1.5mu\mathsf{T}}P_{0})^{\mkern-1.5mu\mathsf{T}}P_{{\alpha^{*}}}+P_{{\alpha^{*}}}(A_{0}-BB^{\mkern-1.5mu\mathsf{T}}P_{0})$
$=-P_{0}A_{{\alpha^{*}}}-A_{{\alpha^{*}}}^{\mkern-1.5mu\mathsf{T}}P_{0}-P_{{\beta^{*}}}A_{{\delta^{*}}}-A_{{\delta^{*}}}^{\mkern-1.5mu\mathsf{T}}P_{{\beta^{*}}}-P_{{\delta^{*}}}A_{{\beta^{*}}}$
$\quad{}-{}A_{{\beta^{*}}}^{\mkern-1.5mu\mathsf{T}}P_{{\delta^{*}}}+P_{{\beta^{*}}}BB^{\mkern-1.5mu\mathsf{T}}P_{{\delta^{*}}}+P_{{\delta^{*}}}BB^{\mkern-1.5mu\mathsf{T}}P_{{\beta^{*}}}.$ (13)
Note that for given $r$ and $p$, there exist exactly $\binom{r+p-1}{p}$ unique such multiindices $\alpha$ with the absolut value of $\rvert\alpha\lvert=p$. For such a series expansion of $P$, the feedback law eq. 8 reads
$u(t)=-B^{\mkern-1.5mu\mathsf{T}}\bigl{(}\sum_{\rvert\alpha\lvert\leq p}\rho^{(\alpha)}P_{\alpha}\bigr{)}v(t)=:u_{P_{|\alpha|\leq p}}(t),$ (14)
which with the precomputed $B^{\mkern-1.5mu\mathsf{T}}P_{\alpha}$ is efficiently evaluated by adapting the values of $\rho$. We note that for $p=0$ this formula reduces to the standard linear quadratic regulator (LQR) feedback with respect to a linearization about $v^{*}=0$.
5 Solving the high-dimensional matrix equations
To solve the occurring high-dimensional linear and nonlinear matrix equations eqs. 11, 12 and 13, some reformulations and state-of-the-art approaches are needed. Assume that the dimensions of the input and output operators $m,\ell$ are much smaller than the state-space dimension $n$ and that $A_{0}$ in eq. 10 is sparsely populated. The first equation eq. 11 is the standard factorized Riccati equation, which under the assumptions above can be efficiently solved for a low-dimensional approximation of the solution $Z_{0}Z_{0}^{\mkern-1.5mu\mathsf{T}}\approx P_{0}$, with $Z_{0}\in\mathbb{R}^{n\times k}$ and $k\ll n$, e.g., via the low-rank Newton-Kleinman iteration [7]. For the two types of Lyapunov equations eqs. 12 and 13, some reformulations are needed to efficiently compute the solutions in the large-scale sparse setting; namely, the right-hand sides need to be reformulated into symmetric low-rank factorizations. In the case of eq. 12, given the approximation $Z_{0}Z_{0}^{\mkern-1.5mu\mathsf{T}}\approx P_{0}$, the right-hand side can be rewritten as
$-A_{\alpha}^{\mkern-1.5mu\mathsf{T}}P_{0}-P_{0}A_{\alpha}\approx-\begin{bmatrix}A_{\alpha}^{\mkern-1.5mu\mathsf{T}}Z_{0}&Z_{0}\end{bmatrix}J\begin{bmatrix}Z_{0}^{\mkern-1.5mu\mathsf{T}}A_{\alpha}\\ Z_{0}^{\mkern-1.5mu\mathsf{T}}\end{bmatrix},$
where $J=\begin{bmatrix}0&I_{k}\\ I_{k}&0\end{bmatrix}$. For Lyapunov equations eq. 12 with such symmetric factorized right-hand sides, the ADI approach has been extended to compute appropriate solution approximations of the form $L_{\alpha}D_{\alpha}L_{\alpha}^{\mkern-1.5mu\mathsf{T}}\approx P_{\alpha}$; see [18]. Similarly, we can derive for the right-hand side of eq. 13 low-rank factorizations of the form $LDL^{\mkern-1.5mu\mathsf{T}}$, where
$L^{\mkern-1.5mu\mathsf{T}}=\begin{bmatrix}Z_{0}^{\mkern-1.5mu\mathsf{T}}A_{{\alpha^{*}}}\\ Z_{0}^{\mkern-1.5mu\mathsf{T}}\\ L_{{\beta^{*}}}^{\mkern-1.5mu\mathsf{T}}(A_{{\delta^{*}}}-BK_{{\delta^{*}}})\\ L_{{\delta^{*}}}^{\mkern-1.5mu\mathsf{T}}(A_{{\beta^{*}}}-BK_{{\beta^{*}}})\\ L_{{\beta^{*}}}^{\mkern-1.5mu\mathsf{T}}\\ L_{{\delta^{*}}}^{\mkern-1.5mu\mathsf{T}}\end{bmatrix},D=\begin{bmatrix}J&0&0\\ 0&0&D_{{\beta^{*}},{\delta^{*}}}\\ 0&D_{{\beta^{*}},{\delta^{*}}}&0\end{bmatrix},$
with $D_{{\beta^{*}},{\delta^{*}}}=\begin{bmatrix}D_{{\beta^{*}}}&0\\ 0&D_{{\delta^{*}}}\end{bmatrix}$, the low-rank factorizations $L_{{\beta^{*}}}D_{{\beta^{*}}}L_{{\beta^{*}}}^{\mkern-1.5mu\mathsf{T}}\approx P_{{\beta^{*}}}$ and $L_{{\delta^{*}}}D_{{\delta^{*}}}L_{{\delta^{*}}}^{\mkern-1.5mu\mathsf{T}}\approx P_{{\delta^{*}}}$, and the feedback matrices $K_{{\beta^{*}}}=B^{\mkern-1.5mu\mathsf{T}}P_{{\beta^{*}}}$ and $K_{{\delta^{*}}}=B^{\mkern-1.5mu\mathsf{T}}P_{{\delta^{*}}}$.
6 Numerical experiments
The code, data and results of the numerical experiments presented here can be found at [15].
2345678reduced dimension $r$12345number of clusters $q$Refer to caption1.0e-042.0e-043.0e-04
Figure 1: Grid search for the selection of $r$ and $q$ with lowest averaged reconstruction error.
Table 1: Model complexity of the different PAE schemes and POD: # stands for number of. Further, we distinguish between linear (lin.) and nonlinear (nonl.) mappings in the en–/decoding.
Hyperparameters \  Scheme POD-5 PAE-1.5 PAE-3.5 PAE-3.5 (p=1)
reduced dimension $r$ 5 5 5 5
#clusters $q$ 1 1 3 3
#encoding parameters 255970 42810 42810 42810
#encoding layers 1(lin.) 16(nonl.) 16(nonl.) 16(nonl.)
#decoding parameters 255970 255970 15+767910 255970
#decoding layers 1 1 1(nonl.)+1(lin.) 1
We consider the same problem, numerical setup, and the same data points for training the models as in [16]. In short, the flow control problem of stabilizing the unstable steady-state of the flow behind a cylinder is modeled through the difference $v$ of the velocity to the steady-stated and observed through outputs $y=Cv$. It is determined by the semi-discrete Navier-Stokes equations
$M\dot{v}=N(v)v+Bu,\quad y=Cv,\quad v(0)=0\in\mathbb{R}^{n},$ (15)
[…]
Figure 2: Snapshots of the flow simulation setup.
We consider the steady-state solution as target state and aim for the stabilization of the difference system eq. 15. In order to trigger instabilities and for the collection of training data, a test input
$u(t)=\begin{bmatrix}\sin(t)&0\end{bmatrix}^{\mkern-1.5mu\mathsf{T}},$ (16)
is applied. As described in [14], the chosen deep convolutional encoder leverages depthwise separable convolutions along with a sparse interpolation matrix to mitigate computational complexity and memory requirements in this rather high-dimensional setup. As a result the model complexity in terms of parameters that define the encoder and decoder can be lower than for, say, a comparable POD reduction; see Table 1. For training the PAEs, we adopt a semi-supervised learning methodology with a combined loss function
$\mathcal{L}(v^{(i)}):=\lambda\|\tilde{v}^{(i)}-v^{(i)}\|_{M}-l(\rho^{(i)})\cdot\log(c(\rho^{(i)})),$
where $v^{(i)}$ is the $i$-th data point, $\rho^{(i)}:=\mu(v^{(i)})$ and $\tilde{v}^{(i)}:=\nu(\rho^{(i)})$ for the corresponding encoder and decoder $\mu$ and $\nu$, where $l$ assigns labels that were precomputed by $k$-means clustering in $\mathbb{R}^{r}$, and where $\lambda>0$ is a weighting factor between the reconstruction error (measured in the $M$-weighted discrete $L^{2}$-norm) and the clustering error. We note that the chosen cross-entropy loss is commonly used to smoothly assign labels and that the labeling $l$ is defined such that the target $\rho^{*}=0$ is assigned to the first cluster. As it is standard practice, the loss is computed as the average over a batch of data points. In our experiments, we consider $401$ data points equally distributed in the time interval $[0,0.5]$ as training data and we use the Adam optimizer with a learning rate of $0.005$, with $3200$ epochs, a batch size of $64$ and the loss weight $\lambda=100$.
6.1 Reconstruction errors
Table 2: Average of the reconstruction errors (scaled with $10^{3}$) of POD-r and PAE-q.r on the time range $[0,0.5]$ used for training.
Scheme \  r 2 3 4 5 6 7 8
POD-r 1.340 0.610 0.305 0.149 0.076 0.038 0.020
PAE-1.r 1.335 0.614 0.326 0.181 0.118 0.121 0.087
PAE-3.r 0.189 0.135 0.104 0.066 0.075 0.097 0.079
PAE-3.r(p=1) 17.670 7.501 5.217 5.628 2.745 6.794 2.307
We evaluate the reconstruction performance of the PAE in the averaged $M$-norm error
$\frac{1}{N}\sum_{i=1}^{N}\|v^{(i)}-\tilde{v}^{(i)}\|_{M},$
where $v^{(i)}$ is the $i$-th data point and $\tilde{v}^{(i)}=\nu(\mu(v^{(i)}))$ for the corresponding encoder and decoder $\mu$ and $\nu$. As laid out in Table 2, the PAE-1.r is comparable to the standard linear approach of POD for low dimensions $r=2,3,4$, whereas the nonlinear encoding through the cluster selection significantly improves the reconstruction in low dimensions. As for the series expansion of the control law, only the first-order approximation is relevant, we tabulate this reconstruction error as well. Note that the bad average performance is due to the strong deviation for larger $t$, i.e. larger values of $\rho$, whereas for small $t$, the first-order approximation closely aligns with the nonlinear approach; see Figure 3.
0.00.10.20.30.40.5time $t$0.00000.0005PAE-1.5PAE-3.5PAE-3.5 (p=1)
Figure 3: Trajectory with pointwise reconstruction errors $\|v(t^{(i)})-\tilde{v}(t^{(i)})\|_{M}$. The improvement of using $3$ clusters is clearly visible as is the initial alignment with the nonlinear approach PAE-3.5 (and the subsequent deviation from it) of the first-order expansion.
Before turning to the controller design, we use a grid search to identify a suitable parameter setup for the reduced-order LPV approximation. As shown in Figure 1, the pair $(q,r)=(4,5)$ achieves the lowest reconstruction error in the considered domain. Nonetheless, we selected $(q,r)=(3,5)$ as a compromise of both accuracy and complexity.
6.2 SDRE approximation through series expansions
We check the approximation of the SDRE feedback along the example trajectory $v$ generated by the test input eq. 16. For that we compute the true SDRE feedback $u_{P}(t)=-B^{\mkern-1.5mu\mathsf{T}}P(v(t))v(t)$, where $P$ solves the SDRE, and compare to $u_{\tilde{P}}$ and $u_{P_{|\alpha|\leq p}}$ that denote the feedback computed through the LPV approximation in eq. 9 (via PAE-q.5, $q=1,3$) and the expansions eq. 14 for $p=0,1,2$, respectively. The data in Figure 4 shows that the LPV approximation works particularly well for $t\leq 0.5$, which is the training regime, and that the first and second-order expansions significantly improve the zero-order approximation, which would be the LQR gain, but with a slight effect of the additional second-order terms.
${10^{-4}}$${10^{-1}}$$q=1,\quad\gamma=10^{-1}$$q=3,\quad\gamma=10^{-1}$0.00.20.40.60.81.0time $t$${10^{-4}}$${10^{-1}}$$q=1,\quad\gamma=10^{2}$0.00.20.40.60.81.0time $t$$q=3,\quad\gamma=10^{2}$ approximation: $\|u_{P}(t)-u_{\tilde{P}}(t)\|$LQR feedback: $\|u_{P}(t)-u_{P_{|\alpha|\leq 0}}(t)\|$first order: $\|u_{P}(t)-u_{P_{|\alpha|\leq 1}}(t)\|$second order: $\|u_{P}(t)-u_{P_{|\alpha|\leq 2}}(t)\|$
Figure 4: Differences to the exact SDRE feedback along a sample trajectory for feedback approximations through the LPV approximation of the nonlinearity by PAE-q.5, for $q=1,3$, and the corresponding zeroth, first and second-order series expansions of the SDRE feedback.
6.3 Feedback performance
To evaluate the performance of the approximations for the controller design, we simulate the closed loop system with varying the two parameters $\gamma\in\{100,10,1,0.1,0.001\}$ that stands for the penalization factor of $u$ in the underlying quadratic cost functional and the startup time $t_{s}>0$ that marks the time after which the test input is turned off and the feedback is applied. In other words, $\gamma$ is a parameter that weighs on the magnitude of the control and $t_{s}$ regulates how far the system is deferred from the target state when the controller is activated. In Figure 5, for different orders $p=0,1,2$ of approximation, we plot the averaged measured feedback magnitude
$\frac{1}{t_{e}}\Big{(}\int_{t_{s}}^{t_{e}}\|B^{\mkern-1.5mu\mathsf{T}}\bigl{(}\sum_{|\alpha|\leq p}\rho(v(t))^{(\alpha)}P_{\alpha}\bigr{)}v(t)\|^{2}\operatorname{d}t\Big{)}^{\frac{1}{2}},$ (17)
that tends to 00 for $t_{e}\to\infty$ if stabilization is achieved, and that is finite if and only if the simulation has not blown up before a finite end time $t_{e}$.
[…]
Figure 5: Performance index eq. 17 map of the linear and nonlinear feedbacks for PAE reduced-order models with $r=5$, $q=1,3$, the SDRE feedback law expansion eq. 14 for $p=0,1,2$ and varying penalization parameter $\gamma$ and startup times $t_{s}$. The darker the color, the smaller (i.e., the better) the performance index. At the blank spaces, the system blew up. The trajectories of the data marked with * are plotted in Figure 6.
[…]
Figure 6: Measured output and feedback input signals for the distinguished data points from Figure 5.
We plot the values of the performance index eq. 17 for $t_{e}=7.5$ and for various approximations to the nonlinear SDRE feedback in Figure 5 and the trajectories of selected cases in Figure 6. As we have observed before, the $p=0$ expansion in eq. 14, i.e., the LQR feedback, is remarkably performant for the optimal choice of the Tikhonov parameter $\gamma$. However, considering higher order expansions with $p=1,2$ together with the option to adapt the underlying approximation scheme, significantly widens the domain of attraction and can ensure reliable performance in particular for larger values of $\gamma$; see, e.g., Figure 6 (top row) for the trajectory data for a setup with $\gamma=10$, where only the second-order expansion achieves stabilization.
7 Conclusion
In the forward simulation model, the proposed nonlinear autoencoder with smooth clustering significantly improves the parametrization quality at very low dimensions. For the controller design, it was established that for second-order expansions of the SDRE feedback law, only the first-order expansions of the reduced-order LPV coefficient are relevant, which reduces the computational effort by large. Concerning the performance of the resulting feedback, the overall picture is diffuse. While the linear LQR approach, for suitable weight parameters, ensured stability also well away from the origin, the nonlinear feedback design based on different approximation schemes, clearly extended the basin of attraction. As the choice of the approximation was solely based on a single open-loop simulation in a time range that did not cover the range considered in the closed-loop simulation, in future work, we expect to gain additional performance by optimizing the approximation in view of feedback control.
Acknowledgments
Jan Heiland and Yongho Kim were supported by the German Research Foundation (DFG) Research Training Group 2297 “Mathematical Complexity Reduction (MathCoRe)”, Magdeburg. Steffen W. R. Werner acknowledges Advanced Research Computing (ARC) at Virginia Tech for providing computational resources and technical support that have contributed to the results reported within this paper.

### v2

@@
-Deep polytopic autoencoders for low-dimensional linear parameter-varying approximations and nonlinear feedback design
+2025-01-23
 Abstract
@@
 For nonlinear PDEs, this is readily applied to low-dimensional linear parameter-varying (LPV) approximations as they have been exploited for efficient nonlinear controller design via series expansions of the solution to the state-dependent Riccati equation.
-In this work, we develop a polytopic autoencoder for control applications and show how it outperforms standard linear approaches in view of LPV approximations of nonlinear systems and how the particular architecture enables higher order series expansions at little extra computational effort.
-We illustrate the properties and potentials of this approach to computational nonlinear controller design for large-scale systems with a thorough numerical study.
+In this work, we develop a polytopic autoencoder for control applications and show how it improves on standard linear approaches in view of LPV approximations of nonlinear systems.
+We discuss how the particular architecture enables exact representation of target states and higher order series expansions of the nonlinear feedback law at little extra computational effort in the online phase and how the linear though high-dimensional and nonstandard Lyapunov equations are efficiently computed during the offline phase.
+In a numerical study, we illustrate the procedure and how this approach can reliably outperform the standard linear-quadratic regulator design.
+keywords:
+Nonlinear systems, nonlinear feedback control, parameter-varying approximations, Riccati equations, Lyapunov equations
 […]
+[…]
+34H05, 37N35, 65F45, 76D55, 93B52
 1 Introduction
-The design of controllers for nonlinear high-dimensional systems is a challenging task and there exists no generally applicable and computationally feasible approach.
+The design of feedback controllers for nonlinear high-dimensional systems is a challenging task and there exists no generally applicable and computationally feasible approach.
 Here, we consider input-affine systems of the form
-$\dot{v}(t)=f(v(t))+Bu(t),\quad y(t)=Cv(t),$ (1)
-where for time $t>0$, $v(t)\in\mathbb{R}^{n}$ denotes the state, $u(t)\in\mathbb{R}^{m}$ and $y(t)\in\mathbb{R}^{\ell}$ denote the system’s input and output, $f\colon\mathbb{R}^{n}\to\mathbb{R}^{n}$ is a possibly nonlinear function, and $B$ and $C$ are linear input and output operators.
-For the control of systems like eq.
-1, methods like backstepping [17], feedback linearization [19, Chap.
-5.3] and sliding mode control [12] require structural assumptions whereas general approaches based on the Hamilton-Jacobi-Bellman (HJB) equations are only feasible for very moderate system sizes; see, e.g., [8].
-The commonly employed state-dependent Riccati equation (SDRE) approximation to the HJB (see, e.g., [10] for a survey) lifts a significant portion of the computational complexity.
-Nonetheless, this approach became available for large-scale systems only in recently due to further simplifications; see [6] for a linear update scheme and [1] for approximations through series expansions.
-In this work, we expand on the series expansion of the SDRE combined with the approximation of the nonlinear system through a low-dimensional linear parameter-varying (LPV) system.
-This allows us to provide a generic algorithm to nonlinear controller design; see our previous works [16, 11] and related discussions in [1, 2].
-In particular, the results in [16] indicate a measurable improvement of the nonlinear feedback design over a related, well-tuned and robust LQR controller but left space for further performance improvements, for example, through
-  1. 1.
-    higher order expansions of the nonlinear feedback or
-  2. 2.
+$\dot{x}(t)=f(x(t))+Bu(t),\quad y(t)=Cx(t),$ (1)
+where for time $t\geq 0$, $x(t)\in\mathbb{R}^{n}$ denotes the state, $u(t)\in\mathbb{R}^{m}$ and $y(t)\in\mathbb{R}^{\ell}$ denote the system’s input and output, $f\colon\mathbb{R}^{n}\to\mathbb{R}^{n}$ is a possibly nonlinear function, and $B\in\mathbb{R}^{n\times m}$ and $C\in\mathbb{R}^{p\times n}$ are linear input and output operators represented as matrices.
+The general idea of feedback controller design is to identify a mapping $\mathcal{K}\colon x(t)\to u(t)$ so that with the choice of $u=\mathcal{K}(x)$, the system’s state $x$ and output $y$ satisfy desired criteria.
+A typical application is feedback stabilization that seeks to achieve $x(t)\to 0$ as $t\to\infty$ or the more general task of set-point control, namely $x(t)\to x^{\ast}$ for some state $x^{\ast}\in\mathbb{R}^{n}$ that fulfills $f(x^{\ast})=0$.
+For the controller design for systems of the form eq.
+1, methods like backstepping [34], feedback linearization [41, Chap.
+5.3], and sliding mode control [22] require structural assumptions, whereas general approaches based on the Hamilton-Jacobi-Bellman (HJB) equations are only feasible for very moderate system sizes or in combination with model reduction; see, e.g., [16] for polynomial expansions, [23] for tensorized approximations, or [3] for an approach that addresses general PDEs based on first reducing the state-space dimension through, e.g., POD.
+Other efforts try to approximate the solution of the HJB equations by deep neural networks; see [20].
+The commonly employed state-dependent Riccati equation (SDRE) approximation to the HJB equations (see, e.g., [18] for a survey) lifts a significant portion of the computational complexity.
+Nonetheless, this approach became available for large-scale systems only recently due to further simplifications; see [13] for a linear update scheme and [1] for approximations through series expansions for systems of a certain parametric structure.
+These structural assumptions have been lifted in our previous work (see [30]) so that, with the help of model reduction, now a large class of nonlinear control systems can be treated with series expansions of the SDRE.
+The underlying idea is the approximation of the nonlinear system eq.
+1 through a very low-dimensional linear parameter-varying (LPV) system so that the series expansion of any nonlinear feedback law can be formulated with respect to a few parameters rather than the full state; see our previous works [30, 21] and related discussions in [1, 2].
+In particular, the results in [30] indicate a measurable improvement of the nonlinear SDRE feedback over a related, well-tuned and robust linear-quadratic regulator (LQR) but left space for further performance improvements, for example, through
+  • (i)
+    higher order expansions of the nonlinear feedback, or
+  • (ii)
     nonlinear parametrizations of the velocity states.
-Both topics are linked to the question if a nonlinear parametrization can provide better low-dimensional approximations with fewer degrees of freedom and thus mitigating the natural limit, which is commonly expressed through the Kolmogorov $n$-width of any linear model order reduction scheme – like proper orthogonal decomposition (POD) which is a linearly optimal projection scheme; [13].
-There have been many efforts and positive results concerning the use of nonlinear model order reduction approaches using, e.g., polynomial manifold maps [9], learning techniques [13], or the selection of local bases [3].
-To consider higher order expansions of a nonlinear feedback law, a very low-dimensional parametrization is crucial.
-Among others, as we will lay out, the numbers of coefficient that need to be precomputed through large-scale matrix equations equals $\sum_{k=0}^{p}\binom{r+k-1}{k}$, where $r$ is the dimension of the parametrization and $p$ is the order of the expansion.
-For example, for $r=10$ and a first-order expansion, the solve of $1+10=11$ matrix equations was required, cp.
-[16], whereas a second or third-order expansion would require $66$ or $286$ solves.
-However, for a parametrization of size $r=5$, only $6$, $21$, or $56$ matrix equation solves are needed for expansions of first, second and third order, respectively.
-In this work, we rely on our recent development of polytopic autoencoders [14] that combines learning approaches with the selection of local bases in terms of polytope vertices.
-The concept of reconstructing the states in a polytope provides certain advantages over the reconstruction in a linear space.
-Generally, the observed states of a dynamical system are described in a bounded set rather than in an unbounded space.
-With the polytopic approach, the reconstruction is confined to a region where the states have been observed.
-This confinement stands against the freedom to explore the full state space, for controller design, however, it can be beneficial if tailored towards target regimes and since the resulting series expressions will be bounded by design.
-The underlying ideas and major findings of this paper can be summarized as follows: We recall in Section 2 that a general nonlinear control system can be approximated by a low-dimensional LPV system, which can be exploited for controller design in particular if the parameter dependency itself is linear.
+For higher order expansions of nonlinear feedback laws, a low-dimensional parametrization of the states is crucial since the number of coefficients grows drastically with the dimension of the parametrization and the order of the expansion so that both topics are linked to the single question if a nonlinear parametrization can provide sufficiently accurate low-dimensional approximations with only few degrees of freedom.
+It is known that linear parametrizations like the (linearly) optimal approach of proper orthogonal decomposition (POD, see, e.g., [26]) have a natural limit on the approximation accuracy, which is commonly expressed through the Kolmogorov $n$-width; see [37].
+Thus, there have been many efforts and positive results concerning the use of nonlinear model order reduction approaches including, for example, polynomial manifold maps [9, 17] or general nonlinear autoencoders [26, 40, 38].
+An alternative for accurate low-dimensional parametrizations is found in the use of local bases [4]; see the recent works that consider local POD bases state reconstruction on subsets of the given dataset [5, 6, 33] or our previous work on local bases approximations via autoencoders [32, 31].
+Basically, local bases are designed for subsets of states considered to have high similarities.
+However, local bases require a mechanism to identify the subsets and which basis is to be used.
+Standard approaches use clustering for this purpose, for example, $k$-means clustering, fuzzy $c$-means clustering, and deep clustering [8, 24, 25].
+In this work, we call on autoencoders with smooth basis selection and present a purely computational approach to design reduced-order parametrizations of nonlinear equations and provide algorithms and basic results for applications in nonlinear feedback control.
+Therefore, we extend our recently developed framework of polytopic autoencoders [32] that combines an autoencoder with the selection of local bases in terms of polytope vertices towards the use in set-point control.
+In particular, we directly impose the set-point condition $f(x^{\ast})=0$ onto the autoencoder so that we obtain a nonlinear encoding of very small dimension with arbitrary accuracy in the relevant part of the state space.
+To this end, we derive explicit expressions for the Jacobian of the decoder, which we then can exploit in theory and practice for analysis and synthesis of higher order expansions of the SDRE feedback.
+The precomputation of the expansion coefficients requires the solution of large-scale matrix-valued algebraic equations of a nonstandard form.
+Here, we call and expand on recent advances in the solution of indefinite matrix Lyapunov and Riccati equations; see [35, 39].
+We note that the derived representation through a polytopic LPV system is suited for controller design on the base of linear matrix inequalities (LMIs); see, e.g., [7].
+However, since the reduction only considers the structure but not the high-dimensional state space, the systems to solve are far too large for state-of-the-art LMI solvers.
+We exemplify the procedure and potentials of this general-purpose approach to nonlinear feedback design with a numerical example based on the two-dimensional flow past a cylinder.
+The overall structure of this paper can be summarized as follows: We recall in Section 2 that a general nonlinear control system can be approximated by a low-dimensional LPV system, which can be exploited for controller design in particular if the parameter dependency itself is linear.
 In view of providing a performant but maximally low-dimensional parametrization, we resort to clustered polytopic autoencoders that encode the states of a dynamical system in a polytope and decode them with locally adapted bases; see Section 3.
 An interesting result in view of controller design is that this nonlinear decoding has a Taylor expansion with vanishing second-order terms (Lemma 1).
-For the actual controller design in Section 4, we consider series expansions of state-dependent Riccati equation as it has been proposed and exercised before.
+For the actual controller design in Section 4, we consider series expansions of state-dependent Riccati equation.
 Here, we extend the formulas to consider nonlinear parameter dependencies and higher order terms in the expansion.
-Finally, we report numerical experiments that illustrate the performance improvement of the polytopic autoencoder over standard approximation schemes, that explore the different levels of approximation to the underlying nonlinear SDRE feedback design, and that investigate the gain in controller performance for a commonly considered flow control problem; see Section 6.
-2 Low-dimensional LPV approximations
+We discuss and showcase, how the relevant large-scale matrix equations can be solved numerically.
+Finally, we report numerical experiments that illustrate the performance improvement of the polytopic autoencoder over standard approximation schemes that explore the different levels of approximation to the underlying nonlinear SDRE feedback design, and that investigate the gain in controller performance for a commonly considered flow control problem; see Section 6.
+As a result, we obtain a model-based nonlinear controller of reduced complexity, with the reduction tailored to snapshot data.
+The paper is concluded in Section 7.
+2 Low-dimensional linear parameter-varying approximations
 For nonlinear systems like eq. 1, we consider quasi-LPV approximations of the form
-$\dot{v}(t)=A\bigl{(}\rho(v(t))\bigr{)}\,v(t)+Bu(t),\quad y(t)=Cv(t).$ (2)
-Such systems can be derived from a state-dependent coefficient representation such as $f(v)=\widetilde{A}(v)v$ and a parametrization of the state $v(t)\approx\tilde{v}(t)=\nu(\mu(v(t))$, where $\mu$ is an encoder and $\nu$ is a decoder.
-In this case, we set $\rho=\mu(v)$ and $A(\rho):=\widetilde{A}(\nu(\rho))$.
-Note that the parameter depends on the state, hence the term quasi-LPV.
+$\dot{x}(t)=A\bigl{(}\rho(x(t))\bigr{)}x(t)+Bu(t),\quad y(t)=Cx(t).$ (2)
+Such systems can be derived from an (exact) state-dependent coefficient representation such as $f(x)=\widetilde{A}(x)x$ and a parametrization of the state $x(t)\approx\tilde{x}(t)=\nu(\mu(x(t))$, where $\mu$ is an encoder and $\nu$ is a decoder; see [30].
+In this case, we set $\rho=\mu(x)$ and $A(\rho):=\widetilde{A}(\nu(\rho))$.
+Note that the parameter $\rho$ depends on the state, hence the term quasi-LPV.
 Remark 1.
-In the example application we consider later, namely flows, we will work with coefficients $\widetilde{A}$ that are linear in the variable so that linear structures in the decoding $\rho\to\tilde{v}$ will lead to linear parameter dependencies as they have been proven beneficial for controller design; see, e.g., [4].
-3 Polytopic parametrizations in clusters for controller design
-As opposed to linear model order reduction approaches, where the state is parametrized in a linear subspace, in polytopic autoencoders (PAEs), the state is represented through coordinates with respect to vertices of an $R$-dimensional polytope.
-As a polytope is bounded, without loss of generality, the vertices $\tilde{v}_{k}$ can be scaled and shifted such that the coordinates $\theta(t)\in\mathbb{R}^{R}$ that represent the reconstruction $\tilde{v}(t)$ in the polytope will be positive and sum up to one, i.e.,
-$\tilde{v}(t)=\sum_{k=1}^{R}\theta_{k}(t)\tilde{v}_{k},\quad\text{with}\leavevmode\nobreak\ \theta(t)\geq 0,\quad\sum_{k=1}^{R}\theta_{k}(t)=1.$
-This design is not capable of exactly parametrizing a target value $v^{*}=0$ unless the support vectors $\tilde{v}_{k}$ come with a specific linear dependency.
-To explicitly enable the reconstruction of the zero target value and to ensure that the series expansion of the SDRE is about $\theta(\tilde{v})=0$, we include $\tilde{v}_{0}=0$ in the reconstruction basis of the training, where we impose the positivity and summation constraint on the $\theta$ variable extended by $\theta_{0}$, i.e.,
-$\theta_{0}(t)+\sum_{k=1}^{R}\theta_{k}(t)=1.$ (3)
-This way, the tuple $(1,\theta(t))=(1,0)\in\mathbb{R}^{R+1}$ will be a feasible value in the extended parameter space that represents the target value.
-Following previous results that showed how clustering, i.e., how the choice of local bases can improve the reconstruction, we combine the polytopic parametrization with smooth clustering networks as we have developed it in [14].
-In short, we will use coordinates $\theta=c(\rho)\otimes\rho$, where $\rho\in\mathbb{R}^{r}$ denotes the actual latent variable that encodes $v$ and where $c$ is a smooth clustering network that blends between $q$ sets of local bases for the reconstruction.
-Taking into consideration a zero target value, we propose a polytopic autoencoder that consists of three components:
-  1. 1.
+In the example application of fluid flow models, the coefficient function $\widetilde{A}$ is readily available; cf.
+[28].
+Here, the function $\widetilde{A}$ is linear so that a linear decoding will lead to linear parameter dependencies as they have been proven beneficial for controller design; see, e.g., [7].
+In what follows, we refer by $\rho_{i}$ to the $i$-th component of a vector-valued quantity $\rho$.
+To reduce the notation, we omit dependencies and sometimes write, e.g., $\rho$ or $\rho(t)$ instead of $\rho(x(t))$.
+Furthermore, we will denote the Kronecker product by $\otimes$.
+3 Polytopic parametrizations for control
+In polytopic autoencoders (PAEs), the state is represented through generalized barycentric coordinates with respect to $R\in\mathbb{N}$ vertices $\{\tilde{x}_{k}\}_{k=1}^{R}\subset\mathbb{R}^{n}$ of a polytope; see, e.g., our previous work [32] for a more detailed introduction.
+As a polytope is bounded, one can always assume without loss of generality that the coordinates $\rho(t)\in\mathbb{R}^{R}$ in the reconstruction will be positive and sum up to one, i.e., for the $\tilde{x}(t)$ in the polytope it holds that
+$\tilde{x}(t)=\sum_{k=1}^{R}\rho_{k}(t)\tilde{x}_{k},\quad\text{with }\rho(t)\geq 0\text{ and }\sum_{k=1}^{R}\rho_{k}(t)=1.$ (3)
+An immediate advantage of the polytopic parametrization for the intended controller design through series expansions results from the boundedness of the reconstruction, which keeps the controller gains bounded so that we can expect good approximations in series expansions while possible scaling issues will be detected when precomputing the expansion coefficients.
+For the application in controller design, the consistency at the target state $x^{\ast}$ is important.
+We note that a design like eq.
+3 is not capable of exactly parametrizing a target value $x^{\ast}=0$ unless the vertices $\tilde{x}_{k}$ come with a specific linear dependency.
+Moreover, for the intended series expansions around the target state, accuracy is guaranteed if $\|\rho(t)\|$ is small, which, however, contradicts the summation condition in (3).
+To address these two requirements, we include $\tilde{x}_{0}=0$ in the training of the reconstruction basis, where we impose the positivity and summation constraint on the $\rho$ variable extended by $\rho_{0}$, i.e.,
+$\rho_{0}(t)+\sum_{k=1}^{R}\rho_{k}(t)=1.$ (4)
+This way, the tuple $(1,\rho(t))=(1,0)\in\mathbb{R}^{R+1}$ will be a feasible value in the extended parameter space that exactly represents the setpoint $x^{\ast}=0\in\mathbb{R}^{n}$.
+Moreover, smoothness in the parameter space, will ensure that $\rho$ is small around the origin.
+As for the use of local bases, we assume that $q$ different polytopes with vertices
+$\{\tilde{x}_{k}^{(j)}\}_{k=1,\dotsc,R;\,j=1,\dotsc,q}\subset\mathbb{R}^{n}$
+are to be used as local bases for the reconstruction.
+Then, a direct implementation would compute both the local coordinates $\rho$ and the index $j$ of the relevant polytope and reconstruct according to the formula
+$\tilde{x}=\begin{bmatrix}\tilde{x}_{1}^{(1)}&\ldots&\tilde{x}_{R}^{(1)}&\tilde{x}_{1}^{(2)}&\ldots&\tilde{x}_{R}^{(2)}&\ldots&\tilde{x}_{1}^{(q)}&\ldots&\tilde{x}_{R}^{(q)}\end{bmatrix}(\alpha\otimes\rho)$ (5)
+where, in this case, $\alpha=e_{j}$ is the $j$-th canonical basis vector.
+Instead of this hard selection of the $j$-th polytope, we rather use a smooth clustering network (e.g., [43, 25]) that blends between the polytopes by choosing $\alpha=c(\rho)\in\mathbb{R}^{q}$, with $c(\rho)\geq 0$ and $\sum_{\ell=1}^{q}c(\rho)_{\ell}=1$.
+We note that the summation condition on $c(\rho)$ allows to consider $c(\rho)\otimes\rho$ as generalized barycentric coordinates in the polytope spanned by all vertices as in eq.
+5; see [32, Lem 3.1].
+Under the preceding considerations, we propose a polytopic autoencoder designed for set-point control towards $x^{\ast}=0$ that consists of the following three components:
+  1. (i)
     A nonlinear encoder $\mu\colon\mathbb{R}^{n}\to\mathbb{R}^{r+1}$ so that
-    $(\rho_{0},\rho)=\mu(v)=:\operatorname{softmax}(\operatorname{F_{e}}(v)+e_{1}),$
-    with a scalar function $\rho_{0}$ and $\rho(t)\in\mathbb{R}^{r}$, where the function $\operatorname{F_{e}}\colon\mathbb{R}^{n}\to\mathbb{R}^{r+1}$ is a trainable nonlinear convolutional neural network with $\operatorname{F_{e}}(0)=0$ and where $e_{1}\in\mathbb{R}^{r+1}$ is the first unit vector.
-Since $\operatorname{F_{e}}(0)=0$ and with the softmax function
-    $\operatorname{softmax}\colon x\to\left(\tfrac{x_{i}\tanh(ax_{i})}{\sum_{j=1}^{r+1}x_{j}\tanh(ax_{j})}\right)_{i}^{r+1},$
-    for a parameter $a>0$ (that we will set to $a=10$ in the experiments), we ensure that the output $(\rho_{0}(t),\rho(t))$ satisfies
-    $(\rho_{0}(t),\rho_{k}(t))\geq 0\leavevmode\nobreak\ \leavevmode\nobreak\ \text{and}\leavevmode\nobreak\ \leavevmode\nobreak\ \rho_{0}(t)+\sum_{k=1}^{r}\rho_{k}(t)=1;$
-    in particular $\mu(0)=(1,0)$.
-  2. 2.
-    A smooth clustering network $c\colon\mathbb{R}^{r}\to\mathbb{R}^{q}$ so that
+    $(\rho_{0},\rho)=\mu(x)=:\operatorname{softmax}(\operatorname{F_{e}}(x)+e_{1}),$
+    with a scalar function $\rho_{0}$ and $\rho(t)\in\mathbb{R}^{r}$, where $e_{1}\in\mathbb{R}^{r+1}$ is the first unit vector and $\operatorname{F_{e}}\colon\mathbb{R}^{n}\to\mathbb{R}^{r+1}$ is a nonlinear convolutional neural network described via
+    $h_{1}$ $=\mathsf{T}x,$
+    $h_{2}$ $=\mathsf{ELU}(\mathsf{Conv}_{1}(h_{1})),$
+[…]
+    $h_{L}$ $=\mathsf{ELU}(\mathsf{Conv}_{L-1}(h_{L-1})),$
+    $\operatorname{F_{e}}(x)$ $=\mathsf{FC}(h_{L}).$
+    Here, $\mathsf{T}$ denotes the interpolation of spatially distributed data onto a pixel grid as it is needed for the convolutions, $\mathsf{ELU}$ is the exponential linear unit (ELU) activation function [19], $\mathsf{FC}$ is a fully connected layer, and $\mathsf{Conv}_{i}$ is a convolutional layer, $i\in\{1,2,\cdots,L-1\}$ for some depth parameter $L\in\mathbb{N}$.
+Each layer is designed without bias terms to ensure $\operatorname{F_{e}}(0)=0$ so that with the softmax function
+    $\operatorname{softmax}(\xi)_{i}=\tfrac{\xi_{i}\tanh(10\xi_{i})}{\sum_{j=1}^{r+1}\xi_{j}\tanh(10\xi_{j})}$
+    for $i=1,\ldots,r+1$ and $\xi=(\xi_{1},\ldots,\xi_{r+1})\in\mathbb{R}^{r+1}$, we achieve, by design, that the output $(\rho_{0}(t),\rho(t))$ satisfies
+    $(\rho_{0}(t),\rho_{k}(t))\geq 0\quad\text{and}\quad\rho_{0}(t)+\sum_{k=1}^{r}\rho_{k}(t)=1;$
+    and in particular $\mu(0)=(1,0)$.
+  2. (ii)
+    A smooth clustering network $c\colon\mathbb{R}^{r}\to\mathbb{R}^{q}$, with
     $c(\rho)=\operatorname{softmax}(\operatorname{F_{c}}(\rho)+\hat{e}_{1}),$ (6)
-    where $\operatorname{F_{c}}$ is a trainable neural network with $\operatorname{F_{c}}(0)=0$ and where $\hat{e}_{1}$ is the first unit vector in $\mathbb{R}^{q}$.
-By design, for the output $c(\rho)$, we have that
+    where $\hat{e}_{1}$ is the first unit vector in $\mathbb{R}^{q}$ and
+    $\operatorname{F_{c}}(\rho)=\mathsf{ELU}(\mathsf{FC}(\rho)),$
+    with no bias, i.e., $\operatorname{F_{c}}(0)=0$. By design, for the output $c(\rho)$, we have that
     $c(\rho(t))\geq 0\quad\text{and}\quad\sum_{i=1}^{q}c(\rho(t))_{i}=1$
-    hold, so that $(\theta_{0},\theta):=(\rho_{0},c(\rho)\otimes\rho)$ satisfies the positivity and summation conditions eq. 3.
-  3. 3.
-    A trainable linear map $W\in\mathbb{R}^{n\times qr}$ that realizes the reconstruction as
-    $\tilde{v}=W\big{(}c(\rho)\otimes\rho\big{)}.$ (5)
-We lay out the design and motivation in the following remarks:
-  • The property that $\operatorname{F_{e}}(0)=0$ and $\operatorname{F_{c}}(0)=0$ is achieved by using (convolutional) nonlinear layers without bias terms and with activation functions that cross the origin.
+    hold, so that $(\rho_{0},c(\rho)\otimes\rho)$ satisfies the positivity and summation conditions eq. 4.
+  3. (iii)
+    A trainable linear map $W\in\mathbb{R}^{n\times qr}$ that realizes the reconstruction (cf. eq. 5) as
+    $\tilde{x}=W\big{(}c(\rho)\otimes\rho\big{)}.$ (7)
+A few comments on the design of the autoencoder and motivation:
+  • Generally, the property that $\operatorname{F_{e}}(0)=0$ and $\operatorname{F_{c}}(0)=0$ can be achieved by using standard (convolutional) linear layers without bias terms and with any activation function that crosses the origin.
   • The use of the extra $\rho_{0}$ variable ensures that $\rho$ is zero at the origin despite being valid coordinates in a polytope.
-  • The design of the clustering network ensures $c(0)=e_{1}$ so that values of $\rho$ close to zero are associated with the first cluster and that $c(\mu(0))\otimes\mu(0)=(1,0)\in\mathbb{R}^{qr}$ represents the target vector $v^{*}=0$ exactly.
-  • With using the Kronecker product $c(\rho)\otimes\rho$, the $k$-th $r$ columns of $W$ denote the supporting basis for reconstruction in the $k$-th cluster; see [14].
+  • The smooth clustering network, as opposed to a standard clustering, also ensures differentiability of the network as it will become important for the series expansion.
+  • The design of the clustering network ensures that $c(0)=e_{1}$ so that values of $\rho$ close to zero are associated with the first cluster and that $c(\mu(0))\otimes\mu(0)=(1,0)\in\mathbb{R}^{qr}$ represents the target vector $x^{\ast}=0$ exactly.
+With the following lemma, we summarize the resulting properties when employing the autoencoder in series expansions of the coefficient function.
 Lemma 1.
-With $\operatorname{softmax}\colon x\to\left(\tfrac{x_{i}\tanh(ax_{i})}{\sum_{j=1}^{q}x_{j}\tanh(ax_{j})}\right)_{i}^{q}$, it holds that the Jacobian $c^{\prime}$ of the clustering network $c$ as defined in eq.
-4 is zero at $\rho=0$.
-In particular, the second-order Taylor expansion of $\rho\to A(\rho)$ coincides with the first-order expansion and reads
+Given the smooth clustering network $c$ as defined in eq. 6 and the corresponding LPV coefficient function
+$\rho\to A(\rho)=\widetilde{A}\Bigr{(}W\,\bigl{(}c(\rho)\otimes\rho)\Bigr{)}$ (8)
+defined through the linear map $\widetilde{A}\colon\mathbb{R}^{n}\to\mathbb{R}^{n\times n}$ and the decoder eq. 7. Then, it holds that
+  • (a)
+    the coefficient function eq. 8 is differentiable,
+  • (b)
+    the Jacobian of the clustering network $c$ is zero at $\rho=0$, i.e.,
+    $c^{\prime}(\rho)\bigl{|}_{\rho=0}=0\in\mathbb{R}^{q,r},$
+    and
+  • (c)
+    the second-order Taylor expansion of eq. 8 coincides with the first-order expansion and reads
     $A(\rho)\approx A(0)+\sum_{j}^{r}\rho_{j}\widetilde{A}(w_{j}),$ (9)
@@
 Proof.
-By direct computations we find $c^{\prime}(\operatorname{F_{c}}(\rho)+\hat{e}_{1})\big{\rvert}_{\rho=0}=c^{\prime}(\hat{e}_{1})=0\in\mathbb{R}^{q\times r}$.
-Since the coefficient map $c(\rho)\otimes\rho\to\widetilde{A}(\tilde{v}(\rho))$ is linear, we can reduce the argument to the nonlinear part $\rho\to c(\rho)\otimes\rho$.
-With $\rho=0$, from
+As for the result (a), we note that, by construction, $c$ is a concatenation of arbitrarily often differentiable functions.
+As for the part (b), by direct computations, we can confirm that $c^{\prime}(\operatorname{F_{c}}(\rho)+\hat{e}_{1})\big{\rvert}_{\rho=0}=c^{\prime}(\hat{e}_{1})=0\in\mathbb{R}^{q\times r}$.
+Since the coefficient map
+$\rho\to\widetilde{A}\Bigr{(}W\,\bigl{(}c(\rho)\otimes\rho)\Bigr{)}$
+is linear in components that concern $W$ and $\widetilde{A}$ (cf. Equation 7 and Remark 1), we can reduce the argument to the nonlinear part $\rho\to c(\rho)\otimes\rho$. With $\rho=0$, from
 $c(\rho+h)\otimes(\rho+h)-c(\rho)\otimes\rho=c(0+h)\otimes h,$
 we derive the first and second-order expansion terms of the map $\rho\to c(\rho)\otimes\rho$ in the increment $h$ at $\rho=0$ as
-$c(0+h)\otimes h=\left(c(0)+c^{\prime}(0)h+o(\lvert h\rvert^{2})\right)\otimes h$
-and as
+$c(0+h)\otimes h=\left(c(0)+c^{\prime}(0)h+\mathcal{R}(h)\right)\otimes h$
+with the remainder term $\mathcal{R}(h)\in o(|h|)$, meaning that $\mathcal{R}$ approaches 00 faster than $h$ itself.
+Thus, the second-order expansion of $\rho\to c(\rho)\otimes\rho$ around the origin 00 reads
 $c(\rho)\otimes\rho\approx c(0)\otimes\rho+(c^{\prime}(0)\rho)\otimes\rho=c(0)\otimes\rho+(c^{\prime}(0)\otimes I)(\rho\otimes\rho),$
 where $c^{\prime}(0)\in\mathbb{R}^{q\times r}$ is the Jacobian of $c$ at $\rho=0$.
-With $c^{\prime}(0)=0$, the second-order terms vanish so that with $c(0)=\hat{e}_{1}$, the first (and second-order) approximation to the polytopic reconstruction reads
-$\tilde{v}(\rho)\approx W\left(\hat{e}_{1}\otimes\rho\right)=\sum_{j=1}^{r}\rho_{j}w_{j}.$ (7)
-Therefore, formula eq. 6 follows from the linearity of $\widetilde{A}$. ∎
+With $c^{\prime}(0)=0$, the second-order terms vanish so that with $c(0)=\hat{e}_{1}$, the second-order expansions reduces to the first order approximation of the polytopic reconstruction that reads
+$\tilde{x}(\rho)\approx W\left(c(0)\otimes\rho\right)=W\left(\hat{e}_{1}\otimes\rho\right)=W\begin{bmatrix}\rho\\ 0\\ \vdots\\ 0\end{bmatrix}=\sum_{j=1}^{r}\rho_{j}w_{j}.$
+Finally, formula eq. 9 follows from the linearity of $\widetilde{A}$ and $W$. ∎
+Remark 2.
+We note the following consistency property in the smooth clustering.
+By construction, the first $r$ columns of $W$ describe the local basis for the reconstruction in the cluster associated with $\rho=0$.
+Accordingly, the linearization around the origin only considers the basis that is associated with the origin.
 4 Higher order series expansions of LPV approximations in SDRE feedback design
-For the quasi-LPV system eq. 2, the nonlinear SDRE feedback design defines the input $u$ as
-$u(t)=-B^{\mkern-1.5mu\mathsf{T}}P(\rho(v(t)))v(t),$ (8)
+Although, as shown in Lemma 1, the second-order expansion of the coefficient function $A$ coincides with the expansion of first order, the feedback law will still exhibit second-order dependencies as we will lay out in this section.
+For the quasi-LPV system eq.
+2, the nonlinear SDRE feedback design defines the input $u$ as
+$u(t)=-B^{\mkern-1.5mu\mathsf{T}}P(\rho(x(t)))x(t),$ (10)
 where $P(\rho)$ solves the SDRE
 $A(\rho)^{\mkern-1.5mu\mathsf{T}}P(\rho)+P(\rho)A(\rho)-P(\rho)BB^{\mkern-1.5mu\mathsf{T}}P(\rho)=-C^{\mkern-1.5mu\mathsf{T}}C.$ (11)
-Since the repeated solve of eq. 9 is costly, series expansions of the feedback law eq. 8 have been proposed.
-We briefly derive the formulas for higher order expansions with nonlinear parameter dependencies in $A$.
+We note that for quasi-LPV systems the parameter changes continuously with the states.
+Since the repeated solve of eq.
+11 is costly, series expansions of the feedback law eq.
+10 have been proposed.
+We now briefly derive the formulas for higher order expansions of the solution to eq.
+11 with nonlinear parameter dependencies in $A$.
 For linear parameter dependencies, the first-order approximation has been developed in [10] for a single parameter dependency and has been extended in [1] to the multivariate case.
 If the state dependency of the SDRE solution is parametrized through $\rho$, the multivariate Taylor expansion of $P$ about $\rho_{0}=0$ up to order $p$ reads
-$P(\rho)\approx P_{0}+\sum_{1\leq\lvert\alpha\rvert\leq p}\rho^{(\alpha)}P_{\alpha},$
-where $\alpha=(\alpha_{1},\ldots,\alpha_{r})\in\mathbb{N}^{r}$ is a multiindex with $\lvert\alpha\rvert:=\sum_{i=1}^{r}\alpha_{i}$, where $\rho^{(\alpha)}:=\rho_{1}^{\alpha_{1}}\rho_{2}^{\alpha_{2}}\cdots\rho_{r}^{\alpha_{r}}$, and where, importantly, $P_{\alpha}$ are constant; see [16] for more details.
-For nonlinear parameter dependencies, we will consider a truncated series expansion of $A$:
+$P(\rho)\approx P_{0}+\sum_{1\leq\lvert\alpha\rvert\leq p}\rho^{(\alpha)}P_{\alpha},$ (12)
+where $\alpha=(\alpha_{1},\ldots,\alpha_{r})\in\mathbb{N}^{r}$ is a multiindex with $\lvert\alpha\rvert:=\sum_{i=1}^{r}\alpha_{i}$, where the coefficients are $\rho^{(\alpha)}:=\rho_{1}^{\alpha_{1}}\rho_{2}^{\alpha_{2}}\cdots\rho_{r}^{\alpha_{r}}$, and where, importantly, all $P_{\alpha}$ are constant symmetric matrices, namely
+$P_{\alpha}=\frac{1}{\alpha_{1}!\alpha_{2}!\cdots\alpha_{r}!}\frac{\partial^{\lvert\alpha\rvert}}{\partial_{\rho_{1}}^{\alpha_{1}}\partial_{\rho_{2}}^{\alpha_{2}}\cdots\partial_{\rho_{r}}^{\alpha_{r}}}P(0).$
+We note that, because the function $P$ is not known, this formula is merely stated to confirm that the $P_{\alpha}$ are constant matrices.
+Practically, in what follows, we will compute the $P_{\alpha}$ directly by equating coefficients in truncated series.
+For the nonlinear parameter dependencies in $A(\rho)$, we will consider a truncated series expansion of $A$:
 $A(\rho)\approx A_{0}+\sum_{1\leq\lvert\alpha\rvert\leq p}\rho^{(\alpha)}A_{\alpha},$ (13)
 where $A_{\alpha}$ are constant matrices; see eq.
-6 that describes the first and second order expansion in the considered case with the structure of the nonlinear decoding as $\rho\to W(c(\rho)\otimes\rho)$.
-We insert the expansions of $A$ and $P$ into the SDRE eq.
+9 that describes the first and second-order expansions as follows: For $p=1$, we can identify all $\lvert\alpha\rvert=p$ with $k=1,\dotsc,r$ and obtain $\rho^{(\alpha)}A_{\alpha}=\rho_{k}\widetilde{A}(w_{k})$, where $w_{k}$ is the $k$-th column of $W$.
+Following Lemma 1, in the considered case with the structure of the nonlinear decoding as $\rho\to W(c(\rho)\otimes\rho)$, the second-order expansion coincides with the first-order one.
+We also note that, if $\widetilde{A}$ is known (see Remark 1), then $A_{0}$ is readily inferred without knowledge of $\rho$ or $W$.
+We insert the expansions of $A$ and $P$, namely eqs.
+12 and 13, into the SDRE eq.
 11 and solve for the coefficients $P_{\alpha}$ by matching the individual powers of $\rho$.
-For $\alpha=0$, we obtain the standard Riccati equation
+For $\alpha=0$, we obtain the classical standard Riccati equation
 $A_{0}^{\mkern-1.5mu\mathsf{T}}P_{0}+P_{0}A_{0}-P_{0}BB^{\mkern-1.5mu\mathsf{T}}P_{0}=-C^{\mkern-1.5mu\mathsf{T}}C,$ (14)
-whereas for $\lvert\alpha\rvert=1$, the coefficients $P_{\alpha}$ are obtained from
-$\begin{split}P_{0}\bigl{(}\sum_{\lvert\alpha\rvert=1}\rho^{(\alpha)}A_{\alpha}\bigr{)}+\bigl{(}\sum_{\lvert\alpha\rvert=1}\rho^{(\alpha)}A_{\alpha}\bigr{)}^{\mkern-1.5mu\mathsf{T}}P_{0}\\ {}-{}\sum_{\lvert\alpha\rvert=1}\rho^{(\alpha)}P_{\alpha}BB^{\mkern-1.5mu\mathsf{T}}P_{0}-P_{0}BB^{\mkern-1.5mu\mathsf{T}}\sum_{\lvert\alpha\rvert=1}\rho^{(\alpha)}P_{\alpha}=0\end{split}$
-for arbitrary values of $\rho$. This amounts for the $P_{\alpha}$ to be given via the Lyapunov equations
-$\left(A_{0}-BB^{\mkern-1.5mu\mathsf{T}}P_{0}\right)^{\mkern-1.5mu\mathsf{T}}P_{\alpha}+P_{\alpha}\left(A_{0}-BB^{\mkern-1.5mu\mathsf{T}}P_{0}\right)$
-$=-A_{\alpha}^{\mkern-1.5mu\mathsf{T}}P_{0}-P_{0}A_{\alpha},$ (12)
-for the $r$ multiindices in $\alpha\in\mathbb{N}^{r}$ for which $\lvert\alpha\rvert=1$; see also [1, 16] where the same formulas have been derived.
-Similarly, we may collect the terms in which $\rho$ appears quadratic.
-Let ${\alpha^{*}}$ be a multiindex with $\lvert{\alpha^{*}}\rvert=2$ and let ${\beta^{*}}$ and ${\delta^{*}}$ be those two multiindices with $\lvert{\beta^{*}}\rvert=\lvert{\delta^{*}}\rvert=1$ that satisfy ${\beta^{*}}+{\delta^{*}}={\alpha^{*}}$.
-Then, the coefficient $P_{{\alpha^{*}}}$ that cancels the corresponding contributions in the equation is defined through the Lyapunov equation
-$(A_{0}-BB^{\mkern-1.5mu\mathsf{T}}P_{0})^{\mkern-1.5mu\mathsf{T}}P_{{\alpha^{*}}}+P_{{\alpha^{*}}}(A_{0}-BB^{\mkern-1.5mu\mathsf{T}}P_{0})$
+as it is used for the design of feedback controllers for linear systems.
+Moving on, for $\lvert\alpha\rvert=1$, the coefficients $P_{\alpha}$ are obtained from requiring that the first-order terms fulfill
+$\begin{split}P_{0}\left(\sum_{\lvert\alpha\rvert=1}\rho^{(\alpha)}A_{\alpha}\right)+\left(\sum_{\lvert\alpha\rvert=1}\rho^{(\alpha)}A_{\alpha}\right)^{\mkern-1.5mu\mathsf{T}}P_{0}-\sum_{\lvert\alpha\rvert=1}\rho^{(\alpha)}P_{\alpha}BB^{\mkern-1.5mu\mathsf{T}}P_{0}-P_{0}BB^{\mkern-1.5mu\mathsf{T}}\sum_{\lvert\alpha\rvert=1}\rho^{(\alpha)}P_{\alpha}=0,\end{split}$
+for arbitrary values of $\rho$. This defines the $P_{\alpha}$ as the solutions to the Lyapunov equations
+$\left(A_{0}-BB^{\mkern-1.5mu\mathsf{T}}P_{0}\right)^{\mkern-1.5mu\mathsf{T}}P_{\alpha}+P_{\alpha}\left(A_{0}-BB^{\mkern-1.5mu\mathsf{T}}P_{0}\right)=-A_{\alpha}^{\mkern-1.5mu\mathsf{T}}P_{0}-P_{0}A_{\alpha},$ (15)
+for the $r$ multiindices in $\alpha\in\mathbb{N}^{r}$ for which $\lvert\alpha\rvert=1$ holds; see also [1, 30] where the same formulas have been derived.
+Similarly, we may obtain the $P_{\alpha}$ by setting the terms in which $\rho$ appears quadratic to zero:
+00 0 $=P_{0}\left(\sum_{\lvert\alpha\rvert=2}\rho^{(\alpha)}A_{\alpha}\right)+\left(\sum_{\lvert\alpha\rvert=1}\rho^{(\alpha)}P_{\alpha}\right)\left(\sum_{\lvert\alpha\rvert=1}\rho^{(\alpha)}A_{\alpha}\right)+\left(\sum_{\lvert\alpha\rvert=2}\rho^{(\alpha)}P_{\alpha}\right)A_{0}+\dotsc$
+$\qquad{}-{}\left(\sum_{\lvert\alpha\rvert=1}\rho^{(\alpha)}P_{\alpha}\right)BB^{\mkern-1.5mu\mathsf{T}}\left(\sum_{\lvert\alpha\rvert=1}\rho^{(\alpha)}P_{\alpha}\right)-\left(\sum_{\lvert\alpha\rvert=2}\rho^{(\alpha)}P_{\alpha}\right)BB^{\mkern-1.5mu\mathsf{T}}P_{0}-\dotsc,$
+[…]
+$(A_{0}-BB^{\mkern-1.5mu\mathsf{T}}P_{0})^{\mkern-1.5mu\mathsf{T}}P_{{\alpha^{*}}}+P_{{\alpha^{*}}}(A_{0}-BB^{\mkern-1.5mu\mathsf{T}}P_{0})$ (16)
 $=-P_{0}A_{{\alpha^{*}}}-A_{{\alpha^{*}}}^{\mkern-1.5mu\mathsf{T}}P_{0}-P_{{\beta^{*}}}A_{{\delta^{*}}}-A_{{\delta^{*}}}^{\mkern-1.5mu\mathsf{T}}P_{{\beta^{*}}}-P_{{\delta^{*}}}A_{{\beta^{*}}}$
-$\quad{}-{}A_{{\beta^{*}}}^{\mkern-1.5mu\mathsf{T}}P_{{\delta^{*}}}+P_{{\beta^{*}}}BB^{\mkern-1.5mu\mathsf{T}}P_{{\delta^{*}}}+P_{{\delta^{*}}}BB^{\mkern-1.5mu\mathsf{T}}P_{{\beta^{*}}}.$ (13)
-Note that for given $r$ and $p$, there exist exactly $\binom{r+p-1}{p}$ unique such multiindices $\alpha$ with the absolut value of $\rvert\alpha\lvert=p$.
-For such a series expansion of $P$, the feedback law eq.
-8 reads
-$u(t)=-B^{\mkern-1.5mu\mathsf{T}}\bigl{(}\sum_{\rvert\alpha\lvert\leq p}\rho^{(\alpha)}P_{\alpha}\bigr{)}v(t)=:u_{P_{|\alpha|\leq p}}(t),$ (14)
+$\quad{}-{}A_{{\beta^{*}}}^{\mkern-1.5mu\mathsf{T}}P_{{\delta^{*}}}+P_{{\beta^{*}}}BB^{\mkern-1.5mu\mathsf{T}}P_{{\delta^{*}}}+P_{{\delta^{*}}}BB^{\mkern-1.5mu\mathsf{T}}P_{{\beta^{*}}}.$
+Note that for given $r$ and $p$, there exist exactly $\binom{r+p-1}{p}$ unique such multiindices $\alpha$ such that $\rvert\alpha\lvert=p$.
+Finally, we note that for such a series expansion of $P$, the feedback law eq. 10 reads
+$u(t)=-B^{\mkern-1.5mu\mathsf{T}}\left(\sum_{\rvert\alpha\lvert\leq p}\rho^{(\alpha)}P_{\alpha}\right)x(t)=:u_{P_{\lvert\alpha\rvert\leq p}}(t),$ (17)
 which with the precomputed $B^{\mkern-1.5mu\mathsf{T}}P_{\alpha}$ is efficiently evaluated by adapting the values of $\rho$.
-We note that for $p=0$ this formula reduces to the standard linear quadratic regulator (LQR) feedback with respect to a linearization about $v^{*}=0$.
+We also note that for $p=0$, this formula reduces to the standard linear quadratic regulator (LQR) feedback with respect to a linearization about $x^{\ast}=0$.
 5 Solving the high-dimensional matrix equations
-To solve the occurring high-dimensional linear and nonlinear matrix equations eqs.
-11, 12 and 13, some reformulations and state-of-the-art approaches are needed.
-Assume that the dimensions of the input and output operators $m,\ell$ are much smaller than the state-space dimension $n$ and that $A_{0}$ in eq.
-10 is sparsely populated.
-The first equation eq.
-11 is the standard factorized Riccati equation, which under the assumptions above can be efficiently solved for a low-dimensional approximation of the solution $Z_{0}Z_{0}^{\mkern-1.5mu\mathsf{T}}\approx P_{0}$, with $Z_{0}\in\mathbb{R}^{n\times k}$ and $k\ll n$, e.g., via the low-rank Newton-Kleinman iteration [7].
-For the two types of Lyapunov equations eqs.
-12 and 13, some reformulations are needed to efficiently compute the solutions in the large-scale sparse setting; namely, the right-hand sides need to be reformulated into symmetric low-rank factorizations.
-In the case of eq.
-12, given the approximation $Z_{0}Z_{0}^{\mkern-1.5mu\mathsf{T}}\approx P_{0}$, the right-hand side can be rewritten as
-$-A_{\alpha}^{\mkern-1.5mu\mathsf{T}}P_{0}-P_{0}A_{\alpha}\approx-\begin{bmatrix}A_{\alpha}^{\mkern-1.5mu\mathsf{T}}Z_{0}&Z_{0}\end{bmatrix}J\begin{bmatrix}Z_{0}^{\mkern-1.5mu\mathsf{T}}A_{\alpha}\\ Z_{0}^{\mkern-1.5mu\mathsf{T}}\end{bmatrix},$
-where $J=\begin{bmatrix}0&I_{k}\\ I_{k}&0\end{bmatrix}$.
+To compute the solutions to the occurring high-dimensional matrix equations eqs.
+14, 15 and 16, some reformulations and state-of-the-art approaches are needed.
+In the considered setup, we can exploit that the ranks of $B$ and $C$ are much smaller than the state-space dimension $n$ and that $A_{0}$ in eq.
+13 is sparsely populated, so that the first equation eq.
+14—a standard Riccati equation—can be efficiently solved for a low-rank approximation of the symmetric positive semi-definite stabilizing solution $Z_{0}Z_{0}^{\mkern-1.5mu\mathsf{T}}\approx P_{0}$, with $Z_{0}\in\mathbb{R}^{n\times k}$ and $k\ll n$; see [11, 42], e.g., via the low-rank Newton-Kleinman iteration [14, 39], the incremental subspace iteration [36] or the RADI method [12, 15].
+However, for the two types of Lyapunov equations eq.
+15 and eq.
+16 occurring through the series expansion, some reformulations are needed to compute the solutions in this large-scale setting; namely, the right-hand sides need to be reformulated into symmetric low-rank factorizations.
+Given the approximation $Z_{0}Z_{0}^{\mkern-1.5mu\mathsf{T}}\approx P_{0}$ of the initial Riccati equation eq.
+14, the right-hand side in eq.
+15 can be rewritten as
+$-A_{\alpha}^{\mkern-1.5mu\mathsf{T}}P_{0}-P_{0}A_{\alpha}\approx-\begin{bmatrix}A_{\alpha}^{\mkern-1.5mu\mathsf{T}}Z_{0}&Z_{0}\end{bmatrix}J\begin{bmatrix}Z_{0}^{\mkern-1.5mu\mathsf{T}}A_{\alpha}\\ Z_{0}^{\mkern-1.5mu\mathsf{T}}\end{bmatrix},$ (18)
+where the symmetric yet indefinite center matrix is given by $J=\begin{bmatrix}0&I_{k}\\ I_{k}&0\end{bmatrix}$, with $I_{k}$ denoting the $k$-dimensional identity matrix and $k$ being the number of columns of $Z_{0}$.
 For Lyapunov equations eq.
-12 with such symmetric factorized right-hand sides, the ADI approach has been extended to compute appropriate solution approximations of the form $L_{\alpha}D_{\alpha}L_{\alpha}^{\mkern-1.5mu\mathsf{T}}\approx P_{\alpha}$; see [18].
-Similarly, we can derive for the right-hand side of eq.
-13 low-rank factorizations of the form $LDL^{\mkern-1.5mu\mathsf{T}}$, where
-$L^{\mkern-1.5mu\mathsf{T}}=\begin{bmatrix}Z_{0}^{\mkern-1.5mu\mathsf{T}}A_{{\alpha^{*}}}\\ Z_{0}^{\mkern-1.5mu\mathsf{T}}\\ L_{{\beta^{*}}}^{\mkern-1.5mu\mathsf{T}}(A_{{\delta^{*}}}-BK_{{\delta^{*}}})\\ L_{{\delta^{*}}}^{\mkern-1.5mu\mathsf{T}}(A_{{\beta^{*}}}-BK_{{\beta^{*}}})\\ L_{{\beta^{*}}}^{\mkern-1.5mu\mathsf{T}}\\ L_{{\delta^{*}}}^{\mkern-1.5mu\mathsf{T}}\end{bmatrix},D=\begin{bmatrix}J&0&0\\ 0&0&D_{{\beta^{*}},{\delta^{*}}}\\ 0&D_{{\beta^{*}},{\delta^{*}}}&0\end{bmatrix},$
-with $D_{{\beta^{*}},{\delta^{*}}}=\begin{bmatrix}D_{{\beta^{*}}}&0\\ 0&D_{{\delta^{*}}}\end{bmatrix}$, the low-rank factorizations $L_{{\beta^{*}}}D_{{\beta^{*}}}L_{{\beta^{*}}}^{\mkern-1.5mu\mathsf{T}}\approx P_{{\beta^{*}}}$ and $L_{{\delta^{*}}}D_{{\delta^{*}}}L_{{\delta^{*}}}^{\mkern-1.5mu\mathsf{T}}\approx P_{{\delta^{*}}}$, and the feedback matrices $K_{{\beta^{*}}}=B^{\mkern-1.5mu\mathsf{T}}P_{{\beta^{*}}}$ and $K_{{\delta^{*}}}=B^{\mkern-1.5mu\mathsf{T}}P_{{\delta^{*}}}$.
+15 with such symmetric factorized right-hand sides eq.
+18, the ADI approach has been extended to compute appropriate solution approximations of the form $L_{\alpha}D_{\alpha}L_{\alpha}^{\mkern-1.5mu\mathsf{T}}\approx P_{\alpha}$ following the format of the right-hand side; see [35].
+Furthermore, while the coefficient matrix in eq.
+15, namely
+$A-BB^{\mkern-1.5mu\mathsf{T}}P_{0}=A-BV^{\mkern-1.5mu\mathsf{T}}$ (19)
+with $V=B^{\mkern-1.5mu\mathsf{T}}P_{0}\in\mathbb{R}^{n\times m}$, is dense and high-dimensional, it is not necessary to assemble this matrix at any point because eq.
+19 is a large-scale sparse matrix with low-rank perturbation.
+Solution methods for large-scale Lyapunov equations only rely on matrix-vector operations or the solution to linear systems with eq.
+19.
+The latter can be efficiently computed using the Sherman-Morrison-Woodbury matrix inversion formula or augmented matrix approach [27] such that only the solution of linear systems with large-scale sparse coefficient matrices and matrix-vector products are needed.
+It is important to observe that the coefficient eq.
+19 is Hurwitz, i.e., all eigenvalues lie in the left open half-plane, such that the established solution methods can be applied.
+Similarly, we can derive for the right-hand side of eq. 16 symmetric indefinite low-rank factorizations of the form $LDL^{\mkern-1.5mu\mathsf{T}}$. Thereby, the factor are
+$L^{\mkern-1.5mu\mathsf{T}}=\begin{bmatrix}Z_{0}^{\mkern-1.5mu\mathsf{T}}A_{{\alpha^{*}}}\\ Z_{0}^{\mkern-1.5mu\mathsf{T}}\\ L_{{\beta^{*}}}^{\mkern-1.5mu\mathsf{T}}(A_{{\delta^{*}}}-BK_{{\delta^{*}}})\\ L_{{\delta^{*}}}^{\mkern-1.5mu\mathsf{T}}(A_{{\beta^{*}}}-BK_{{\beta^{*}}})\\ L_{{\beta^{*}}}^{\mkern-1.5mu\mathsf{T}}\\ L_{{\delta^{*}}}^{\mkern-1.5mu\mathsf{T}}\end{bmatrix},$ (20)
+and
+$D=\begin{bmatrix}J&0&0&0&0\\ 0&0&0&D_{{\beta^{*}}}&0\\ 0&0&0&0&D_{{\delta^{*}}}\\ 0&D_{{\beta^{*}}}&0&0&0\\ 0&0&D_{{\delta^{*}}}&0&0\end{bmatrix},$ (21)
+where we have the approximations of the solutions to the Lyapunov equations eq.
+15 as $L_{{\beta^{*}}}D_{{\beta^{*}}}L_{{\beta^{*}}}^{\mkern-1.5mu\mathsf{T}}\approx P_{{\beta^{*}}}$ and $L_{{\delta^{*}}}D_{{\delta^{*}}}L_{{\delta^{*}}}^{\mkern-1.5mu\mathsf{T}}\approx P_{{\delta^{*}}}$, and the intermediate feedback matrices $K_{{\beta^{*}}}=B^{\mkern-1.5mu\mathsf{T}}P_{{\beta^{*}}}$ and $K_{{\delta^{*}}}=B^{\mkern-1.5mu\mathsf{T}}P_{{\delta^{*}}}$.
+As for eq.
+15, this reformulation of the constant terms of eq.
+16 allows the application of the same computational procedures for the solution, for example, the $LDL^{\mkern-1.5mu\mathsf{T}}$-factorized ADI method [35].
+The coefficient matrix in eq.
+16 is the same as in eq.
+15 such that it has the same properties being sparse plus low-rank and Hurwitz.
+Therefore, we can apply the same numerical approaches to solve eq.
+16 as for eq.
+15.
+In both cases of Lyapunov equations eqs.
+15 and 16, we have that the inner dimensions of the reformulated right-hand side terms eqs.
+18, 20 and 21 depend on dimensions of the factors of the solution approximations to eqs.
+14 and 15.
+Therefore, we can expect the dimensions of the right-hand side factors to be larger than the dimensions of the input and output matrices used in the Riccati equation eq.
+14.
+In general, we strongly recommend to compress the right-hand side factors at least to their numerical rank using economy-size QR and eigenvalue decompositions.
+All right-hand sides of Lyapunov equations considered in this paper can be written as indefinite symmetric factorizations of the form $LDL^{\mkern-1.5mu\mathsf{T}}$, with $L\in\mathbb{R}^{n\times k}$ and $D\in\mathbb{R}^{k\times k}$ symmetric.
+For the compression, first, compute the economy-size QR of $L=QR$, with $Q\in\mathbb{R}^{n\times k}$ and $R\in\mathbb{R}^{k\times k}$.
+Next, compute the eigenvalue decomposition of the symmetric matrix
+$RDR^{\mkern-1.5mu\mathsf{T}}=U\Lambda U^{\mkern-1.5mu\mathsf{T}},$ (22)
+where $\Lambda\in\mathbb{R}^{k\times k}$ is a diagonal matrix of real eigenvalues and $U\in\mathbb{R}^{k\times k}$ is the matrix of the corresponding orthogonal eigenvectors.
+Next, the numerical rank is determined based on the diagonal matrix $\Lambda$.
+A typical choice here is to truncate all eigenvalues $\lambda_{i}$ that are small in absolute value, e.g.,
+$\lvert\lambda_{i}\rvert<k\cdot\max_{i=1,\dots,k}\lvert\lambda_{i}\rvert\cdot\epsilon,$
+where $\epsilon$ is the machine precision.
+Note that the original matrix $D$ is assumed to be indefinite such that large positive as well as negative eigenvalues may occur, which need to be kept in the compression.
+For simplicity, assume that the eigenvalues in $\Lambda$ are ordered by absolute size such that we have
+$\Lambda=\begin{bmatrix}\Lambda_{1}&0\\ 0&\Lambda_{2}\end{bmatrix}\quad\text{and}\quad U=\begin{bmatrix}U_{1}&U_{2}\end{bmatrix},$
+where $\Lambda_{1},U_{1}$ contain the eigenvalues and eigenvectors that remain and $\Lambda_{2},U_{2}$ contain the eigenvalues and eigenvectors that are truncated. This results in the compression
+$LDL^{\mkern-1.5mu\mathsf{T}}\approx QU_{1}\Lambda_{1}U_{1}^{\mkern-1.5mu\mathsf{T}}Q^{\mkern-1.5mu\mathsf{T}}=\widetilde{L}\widetilde{D}\widetilde{L}^{\mkern-1.5mu\mathsf{T}},$
+where $\widetilde{L}=QU_{1}$ and $\widetilde{D}=\Lambda_{1}$, which is used to compute the solution to the matrix equations eqs. 15 and 16 instead of the original right-hand sides.
 6 Numerical experiments
 The code, data and results of the numerical experiments presented here can be found at [29].
-2345678reduced dimension $r$12345number of clusters $q$Refer to caption1.0e-042.0e-043.0e-04
-Figure 1: Grid search for the selection of $r$ and $q$ with lowest averaged reconstruction error.
-Table 1: Model complexity of the different PAE schemes and POD: # stands for number of. Further, we distinguish between linear (lin.) and nonlinear (nonl.) mappings in the en–/decoding.
+6.1 Numerical setup
+Figure 1: Spatial domain and snapshots of the flow behind a cylinder in the unstable steady state (left) and in the periodic vortex shedding regime (right); cf. [30].
+For comparability reasons, we consider here the same problem, numerical setup, and the same data points for training the models as in [30].
+In short, the flow control problem of stabilizing the unstable steady-state of the flow behind a cylinder is modeled through the difference of the velocity $v$ to the steady-stated and observed through outputs $y=Cx$.
+It is determined by the semi-discrete Navier-Stokes equations
+$M\dot{x}=N(x)x+Bu,\quad y=Cx,\quad x(0)=0\in\mathbb{R}^{n},$ (23)
+[…]
+Table 1: Model complexity of the different PAE schemes and POD: # stands for number of. We distinguish between linear (lin.) and nonlinear (nonl.) mappings in the encoding and decoding.
 Hyperparameters \  Scheme POD-5 PAE-1.5 PAE-3.5 PAE-3.5 (p=1)
@@
 #decoding layers 1 1 1(nonl.)+1(lin.) 1
-We consider the same problem, numerical setup, and the same data points for training the models as in [16].
-In short, the flow control problem of stabilizing the unstable steady-state of the flow behind a cylinder is modeled through the difference $v$ of the velocity to the steady-stated and observed through outputs $y=Cv$.
-It is determined by the semi-discrete Navier-Stokes equations
-$M\dot{v}=N(v)v+Bu,\quad y=Cv,\quad v(0)=0\in\mathbb{R}^{n},$ (15)
-[…]
-Figure 2: Snapshots of the flow simulation setup.
-We consider the steady-state solution as target state and aim for the stabilization of the difference system eq.
-15.
-In order to trigger instabilities and for the collection of training data, a test input
+We consider the steady-state solution as target state and aim for the stabilization of the difference system eq. 23. To trigger instabilities and to collect training data, a test input
 $u(t)=\begin{bmatrix}\sin(t)&0\end{bmatrix}^{\mkern-1.5mu\mathsf{T}},$ (24)
 is applied.
-As described in [14], the chosen deep convolutional encoder leverages depthwise separable convolutions along with a sparse interpolation matrix to mitigate computational complexity and memory requirements in this rather high-dimensional setup.
+We choose a deep convolutional encoder with depthwise separable convolutions to cope with the high computational complexity and memory requirements in this rather high-dimensional setup; see [32] for details.
 As a result the model complexity in terms of parameters that define the encoder and decoder can be lower than for, say, a comparable POD reduction; see Table 1.
-For training the PAEs, we adopt a semi-supervised learning methodology with a combined loss function
-$\mathcal{L}(v^{(i)}):=\lambda\|\tilde{v}^{(i)}-v^{(i)}\|_{M}-l(\rho^{(i)})\cdot\log(c(\rho^{(i)})),$
-where $v^{(i)}$ is the $i$-th data point, $\rho^{(i)}:=\mu(v^{(i)})$ and $\tilde{v}^{(i)}:=\nu(\rho^{(i)})$ for the corresponding encoder and decoder $\mu$ and $\nu$, where $l$ assigns labels that were precomputed by $k$-means clustering in $\mathbb{R}^{r}$, and where $\lambda>0$ is a weighting factor between the reconstruction error (measured in the $M$-weighted discrete $L^{2}$-norm) and the clustering error.
-We note that the chosen cross-entropy loss is commonly used to smoothly assign labels and that the labeling $l$ is defined such that the target $\rho^{*}=0$ is assigned to the first cluster.
+For training the PAEs, we adopt a semi-supervised learning methodology with the loss function
+$\mathcal{L}(v^{(i)}):=\lambda\|\tilde{x}^{(i)}-x^{(i)}\|_{M}-\ell(\rho^{(i)})\cdot\log(c(\rho^{(i)})),$
+where $x^{(i)}$ is the $i$-th data point, $\rho^{(i)}:=\mu(x^{(i)})$ and $\tilde{x}^{(i)}:=\nu(\rho^{(i)})$ for the corresponding encoder and decoder $\mu$ and $\nu$, where $\ell$ assigns labels that were precomputed by $k$-means clustering in $\mathbb{R}^{r}$, and where $\lambda>0$ is a weighting factor between the reconstruction error (measured in the $M$-weighted $L^{2}$-norm) and the clustering error.
+We note that the chosen cross-entropy loss is commonly used to smoothly assign labels and that the labeling $\ell$ is defined such that the target $\rho^{\ast}=0$ is assigned to the first cluster.
 As it is standard practice, the loss is computed as the average over a batch of data points.
-In our experiments, we consider $401$ data points equally distributed in the time interval $[0,0.5]$ as training data and we use the Adam optimizer with a learning rate of $0.005$, with $3200$ epochs, a batch size of $64$ and the loss weight $\lambda=100$.
+Here, we consider $401$ data points equally distributed in the time interval $[0,0.5]$ as training data, and we use the Adam optimizer with learning rate $0.005$, $3200$ epochs, a batch size of $64$ and the loss weight $\lambda=100$.
 6.2 Reconstruction errors
@@
 PAE-3.r(p=1) 17.670 7.501 5.217 5.628 2.745 6.794 2.307
+0.00.10.20.30.40.5time $t$0.00000.0005PAE-1.5PAE-3.5PAE-3.5 (p=1)
+Figure 2: Trajectory with pointwise reconstruction errors $\|x(t^{(i)})-\tilde{x}(t^{(i)})\|_{M}$.
+The improvement of using $3$ clusters is clearly visible as is the initial alignment with the nonlinear approach PAE-3.5 (and the subsequent deviation from it) of the first-order expansion.
 We evaluate the reconstruction performance of the PAE in the averaged $M$-norm error
-$\frac{1}{N}\sum_{i=1}^{N}\|v^{(i)}-\tilde{v}^{(i)}\|_{M},$
-where $v^{(i)}$ is the $i$-th data point and $\tilde{v}^{(i)}=\nu(\mu(v^{(i)}))$ for the corresponding encoder and decoder $\mu$ and $\nu$.
+$\frac{1}{N}\sum_{i=1}^{N}\|x^{(i)}-\tilde{x}^{(i)}\|_{M},$
+where $x^{(i)}$ is the $i$-th data point and $\tilde{x}^{(i)}=\nu(\mu(x^{(i)}))$ for the corresponding encoder and decoder $\mu$ and $\nu$.
 As laid out in Table 2, the PAE-1.r is comparable to the standard linear approach of POD for low dimensions $r=2,3,4$, whereas the nonlinear encoding through the cluster selection significantly improves the reconstruction in low dimensions.
 As for the series expansion of the control law, only the first-order approximation is relevant, we tabulate this reconstruction error as well.
-Note that the bad average performance is due to the strong deviation for larger $t$, i.e.
-larger values of $\rho$, whereas for small $t$, the first-order approximation closely aligns with the nonlinear approach; see Figure 3.
-0.00.10.20.30.40.5time $t$0.00000.0005PAE-1.5PAE-3.5PAE-3.5 (p=1)
-Figure 3: Trajectory with pointwise reconstruction errors $\|v(t^{(i)})-\tilde{v}(t^{(i)})\|_{M}$.
-The improvement of using $3$ clusters is clearly visible as is the initial alignment with the nonlinear approach PAE-3.5 (and the subsequent deviation from it) of the first-order expansion.
+Note that the bad average performance is due to the strong deviation for larger $t$, i.e., larger values of $\rho$, whereas for small $t$, the first-order approximation closely aligns with the nonlinear approach; see Figure 2.
+2345678reduced dimension $r$12345number of clusters $q$Refer to caption1.0e-042.0e-043.0e-04
+Figure 3: Grid search for the selection of $r$ and $q$ with lowest averaged reconstruction error. A trade-off between the number of clusters and the reduced dimension becomes clearly visible.
 Before turning to the controller design, we use a grid search to identify a suitable parameter setup for the reduced-order LPV approximation.
@@
 6.3 SDRE approximation through series expansions
-We check the approximation of the SDRE feedback along the example trajectory $v$ generated by the test input eq.
-16.
-For that we compute the true SDRE feedback $u_{P}(t)=-B^{\mkern-1.5mu\mathsf{T}}P(v(t))v(t)$, where $P$ solves the SDRE, and compare to $u_{\tilde{P}}$ and $u_{P_{|\alpha|\leq p}}$ that denote the feedback computed through the LPV approximation in eq.
+${10^{-4}}$${10^{-1}}$$q=1,\quad\gamma=10^{-1}$$q=3,\quad\gamma=10^{-1}$0.00.20.40.60.81.0time $t$${10^{-4}}$${10^{-1}}$$q=1,\quad\gamma=10^{2}$0.00.20.40.60.81.0time $t$$q=3,\quad\gamma=10^{2}$ approximation: $\|u_{P}(t)-u_{\tilde{P}}(t)\|$LQR feedback: $\|u_{P}(t)-u_{P_{|\alpha|\leq 0}}(t)\|$first order: $\|u_{P}(t)-u_{P_{|\alpha|\leq 1}}(t)\|$second order: $\|u_{P}(t)-u_{P_{|\alpha|\leq 2}}(t)\|$
+Figure 4: Differences to the exact SDRE feedback along a sample trajectory for feedback approximations through the LPV approximation of the nonlinearity by PAE-q.5, for $q=1,3$, and the corresponding zeroth, first and second-order series expansions of the SDRE feedback.
+The LPV approximation provides the most accurate approximation to the true SDRE feedback in its training regime up to $0.5$.
+The first and second-order approximations also strongly improve over the classical LQR feedback.
+We check the approximation of the SDRE feedback along the example trajectory $x$ generated by the test input eq. 24. For that we compute the true SDRE feedback
+$u_{P}(t)=-B^{\mkern-1.5mu\mathsf{T}}P(x(t))x(t),$
+where $P$ solves the SDRE, and compare to $u_{\tilde{P}}$ and $u_{P_{|\alpha|\leq p}}$ that denote the feedback computed through the LPV approximation in eq.
 11 (via PAE-q.5, $q=1,3$) and the expansions eq.
@@
 The data in Figure 4 shows that the LPV approximation works particularly well for $t\leq 0.5$, which is the training regime, and that the first and second-order expansions significantly improve the zero-order approximation, which would be the LQR gain, but with a slight effect of the additional second-order terms.
-${10^{-4}}$${10^{-1}}$$q=1,\quad\gamma=10^{-1}$$q=3,\quad\gamma=10^{-1}$0.00.20.40.60.81.0time $t$${10^{-4}}$${10^{-1}}$$q=1,\quad\gamma=10^{2}$0.00.20.40.60.81.0time $t$$q=3,\quad\gamma=10^{2}$ approximation: $\|u_{P}(t)-u_{\tilde{P}}(t)\|$LQR feedback: $\|u_{P}(t)-u_{P_{|\alpha|\leq 0}}(t)\|$first order: $\|u_{P}(t)-u_{P_{|\alpha|\leq 1}}(t)\|$second order: $\|u_{P}(t)-u_{P_{|\alpha|\leq 2}}(t)\|$
-Figure 4: Differences to the exact SDRE feedback along a sample trajectory for feedback approximations through the LPV approximation of the nonlinearity by PAE-q.5, for $q=1,3$, and the corresponding zeroth, first and second-order series expansions of the SDRE feedback.
 6.4 Feedback performance
-To evaluate the performance of the approximations for the controller design, we simulate the closed loop system with varying the two parameters $\gamma\in\{100,10,1,0.1,0.001\}$ that stands for the penalization factor of $u$ in the underlying quadratic cost functional and the startup time $t_{s}>0$ that marks the time after which the test input is turned off and the feedback is applied.
-In other words, $\gamma$ is a parameter that weighs on the magnitude of the control and $t_{s}$ regulates how far the system is deferred from the target state when the controller is activated.
-In Figure 5, for different orders $p=0,1,2$ of approximation, we plot the averaged measured feedback magnitude
-$\frac{1}{t_{e}}\Big{(}\int_{t_{s}}^{t_{e}}\|B^{\mkern-1.5mu\mathsf{T}}\bigl{(}\sum_{|\alpha|\leq p}\rho(v(t))^{(\alpha)}P_{\alpha}\bigr{)}v(t)\|^{2}\operatorname{d}t\Big{)}^{\frac{1}{2}},$ (17)
-that tends to 00 for $t_{e}\to\infty$ if stabilization is achieved, and that is finite if and only if the simulation has not blown up before a finite end time $t_{e}$.
 […]
@@
 Figure 6: Measured output and feedback input signals for the distinguished data points from Figure 5.
+The second-order LPV approximation is capable of stabilizing the nonlinear system in situations, in which the first-order approximation begins to fail.
+To evaluate the performance of the approximations for the controller design, we simulate the closed loop system with varying the two parameters $\gamma\in\{100,10,1,0.1,0.001\}$ that stands for the penalization factor of $u$ in the underlying quadratic cost functional and the startup time $t_{s}>0$ that marks the time after which the test input is turned off and the feedback is applied.
+In other words, $\gamma$ is a parameter that weighs on the magnitude of the control and $t_{s}$ regulates how far the system is deferred from the target state when the controller is activated.
+In Figure 5, for different orders $p=0,1,2$ of approximation, we plot the averaged measured feedback magnitude
+$\frac{1}{t_{\operatorname{e}}}\Big{(}\int_{t_{s}}^{t_{\operatorname{e}}}\|B^{\mkern-1.5mu\mathsf{T}}\bigl{(}\sum_{|\alpha|\leq p}\rho(x(t))^{(\alpha)}P_{\alpha}\bigr{)}x(t)\|^{2}\operatorname{d}t\Big{)}^{\frac{1}{2}},$ (25)
+that tends to 00 for $t_{\operatorname{e}}\to\infty$ if stabilization is achieved, and that is finite if and only if the simulation has not blown up before a finite end time $t_{\operatorname{e}}$.
 We plot the values of the performance index eq.
-17 for $t_{e}=7.5$ and for various approximations to the nonlinear SDRE feedback in Figure 5 and the trajectories of selected cases in Figure 6.
+25 for $t_{\operatorname{e}}=7.5$ and for various approximations to the nonlinear SDRE feedback in Figure 5 and the trajectories of selected cases in Figure 6.
 As we have observed before, the $p=0$ expansion in eq.
-14, i.e., the LQR feedback, is remarkably performant for the optimal choice of the Tikhonov parameter $\gamma$.
+17, i.e., the LQR feedback, is remarkably performant for the optimal choice of the Tikhonov parameter $\gamma$.
 However, considering higher order expansions with $p=1,2$ together with the option to adapt the underlying approximation scheme, significantly widens the domain of attraction and can ensure reliable performance in particular for larger values of $\gamma$; see, e.g., Figure 6 (top row) for the trajectory data for a setup with $\gamma=10$, where only the second-order expansion achieves stabilization.
+The reason might be that a larger $\gamma$ seeks for small (in magnitude) control actions, which makes better adaptation pay off while it reduces overshoots that may destabilize the nonlinear controller.
 7 Conclusion
-In the forward simulation model, the proposed nonlinear autoencoder with smooth clustering significantly improves the parametrization quality at very low dimensions.
-For the controller design, it was established that for second-order expansions of the SDRE feedback law, only the first-order expansions of the reduced-order LPV coefficient are relevant, which reduces the computational effort by large.
-Concerning the performance of the resulting feedback, the overall picture is diffuse.
-While the linear LQR approach, for suitable weight parameters, ensured stability also well away from the origin, the nonlinear feedback design based on different approximation schemes, clearly extended the basin of attraction.
-As the choice of the approximation was solely based on a single open-loop simulation in a time range that did not cover the range considered in the closed-loop simulation, in future work, we expect to gain additional performance by optimizing the approximation in view of feedback control.
+In this work, we considered the use of deep polytopic autoencoders and state-dependent Riccati equations for the design of nonlinear controllers.
+We use the concept of linear parameter-varying approximations of nonlinear dynamical systems that are parametrized by deep polytopic autoencoders, and we show that by the particular choice of autoencoders, the parametrization is differentiable and has vanishing second-order terms.
+For the nonlinear feedback design, we extend the series expansion of state-dependent Riccati equations to second-order terms for which we provide computable expressions in form of indefinite Lyapunov and Riccati equations.
+Finally, we test the proposed framework on the classical control example modeling the flow past a cylinder.
+The proposed nonlinear autoencoder with smooth clustering significantly improves the parametrization quality at very low dimensions.
+For controller design, we established that for second-order expansions of the state-dependent feedback law, only the first-order expansions of the reduced-order linear parameter-varying coefficient are relevant, which reduces the computational effort by large.
+Regarding the performance of the resulting feedback, the overall picture is diffuse.
+While the linear quadratic regulator approach, for suitable weight parameters, ensured stability also well away from the origin, the nonlinear feedback design based on different approximation schemes, clearly extended the basin of attraction.
+As the choice of the approximation was solely based on a single open-loop simulation in a time range that did not cover the full range of the closed-loop simulation, in future work, we expect to gain additional performance by optimizing the approximation in view of feedback control.
 Acknowledgments
+Parts of this work have been carried out while Jan Heiland was with the Max Planck Institute for Complex Technical Systems, Magdeburg, and with the Otto von Guericke University Magdeburg.
 Jan Heiland and Yongho Kim were supported by the German Research Foundation (DFG) Research Training Group 2297 “Mathematical Complexity Reduction (MathCoRe)”, Magdeburg.